### v1

# Thompson’s Conjecture for Large Unitary Groups

Abstract.
We prove that every simple unitary group $S$ of sufficiently large order satisfies Thompson’s conjecture, that is, $S$ contains a conjugacy class $C$ such that $C^{2}=S$. The proof relies on exponential character bounds recently obtained in [7] and [15], and a refinement that gives a sharp asymptotic for character ratios in Lie-type groups of type $A$.
The first author was partially supported by the NSF grant DMS-2001349. The second author gratefully acknowledges the support of the NSF (grant DMS-2200850) and the Joshua Barlaz Chair in Mathematics.
1. Introduction
Thompson’s Conjecture asserts that every finite non-abelian simple group $S$ contains a conjugacy class $C$ such that $C^{2}=S$. As far as we know, it first appeared in print in 1985, attributed to Thompson [1]. It is a refinement of Ore’s Conjecture, dating back to 1951 and settled by Liebeck, O’Brien, Shalev, and Tiep [18], which asserts that every element in $S$ is a commutator, or, equivalently, that $\bigcup_{C}CC^{-1}=S$. Thompson’s conjecture says that some single class satisfies $C^{2}=CC^{-1}=S$.
In a series of papers culminating with [3], Ellers and Gordeev proved that Thompson’s conjecture is true except possibly when $S$ is a group of Lie type $X(q)$, where $q<q_{0}$, and $q_{0}$ depends on the type $X$ but is never greater than $9$. In particular, $q_{0}=2$ for split type A, meaning that there are no exceptions of this type. Combining this with methods which work best in high rank, it seems there may be some hope of proving that Thompson’s conjecture holds for all sufficiently large finite simple groups. It suffices to consider the finite simple groups of classical type, and we may take their ranks to be as large as we wish. It is not necessary to consider the family of groups $\mathrm{PSL}_{N}(q)$, just the unitary, symplectic, and orthogonal cases.
In a recent paper [15], we showed there are at most finitely many symplectic exceptions to Thompson’s conjecture. In this paper, we deal with the unitary case.
Theorem 1.
All sufficiently large finite simple groups of the form $S=\mathrm{PSU}_{N}(q)$ satisfy Thompson’s conjecture.
Our basic tool is the Frobenius formula, which implies that if $G$ is any finite group, $C\subset G$ is a conjugacy class, $x\in C$, and $g\in G$, then $g\in C^{2}$ if and only if
(1.1) $\sum_{\chi\in\mathrm{Irr}(G)}\frac{\chi(x)^{2}\bar{\chi}(g)}{\chi(1)}\neq 0,$
where $\mathrm{Irr}(G)$ denotes the set of irreducible characters of $G$. Our basic tool for analyzing character values is Deligne-Lusztig theory.
Following work of Malle-Saxl-Weigel [19] and Larsen-Shalev-Tiep [13], Guralnick and Malle [8, Theorem 1.4] solved the closely related but substantially easier problem of finding, for each $S$, a pair of conjugacy classes, $C_{1}$ and $C_{2}$, such that $C_{1}C_{2}\cup\{1\}=S$. In the setting of [8], it is always possible to find pairs of elements $x_{1}$ and $x_{2}$ such that the set of $\chi$ with $\chi(x_{1})\chi(x_{2})\neq 0$ is very small, often consisting just of the trivial character and the Steinberg character, so (1.1) gives the desired result without much need for delicate estimates of character values. The main obstacle to using (1.1) to prove Thompson’s conjecture is that for any single $x$, there are many characters $\chi$ such that $\chi(x)\neq 0$.
A preview of the approach in this paper is given in [14, Theorem 10.7], where together with Jay Taylor, we proved that $C^{2}\cup\{1\}=S$, when $S:=\mathrm{PSL}_{N}(q)$ is sufficiently large, $N$ is prime, and $C$ is the conjugacy class of an element of $S$ represented by $x\in\mathrm{SL}_{N}(q)$ of order $\frac{q^{N}-1}{q-1}$. This result is harder than finding distinct conjugacy classes whose product covers all non-trivial elements, but it is technically easier than what we do here for several reasons: $\mathrm{PSL}_{N}(q)$ is easier to deal with than $\mathrm{PSU}_{N}(q)$; we do not insist that $1\in C^{2}$, so we are not limited to real classes $C$; and we are just dealing with the case $N$ is prime, so the description of $\{\chi\in\mathrm{Irr}(\mathrm{SL}_{N}(q))\mid\chi(x)\neq 0\}$, for $x$ of order $\frac{q^{N}-1}{q-1}$, is particularly simple: just the cuspidal and unipotent characters.
It turns out to be more convenient to work with $\mathrm{SU}_{N}(q)$, in fact with $\mathrm{GU}_{N}(q)$, than $\mathrm{PSU}_{N}(q)$. The notion of support $\mathsf{supp}(g)$ of any element $g\in\mathrm{GL}_{N}(\bar{\mathbb{F}}_{q})$ was introduced in [13, Definition 4.1.1]; $N-\mathsf{supp}(g)$ is the maximum dimension of any eigenspace of $g$. If $g\in\mathrm{GL}^{\epsilon}(A)=\mathrm{GL}_{N}^{\epsilon}(q)$, where $A=\mathbb{F}_{q}^{N}$ and $\mathrm{GL}^{\epsilon}(A)=\mathrm{GL}(A)=\mathrm{GL}_{N}(q)$ when $\epsilon=+$, and $A=\mathbb{F}_{q^{2}}^{N}$ and $\mathrm{GL}^{\epsilon}(A)=\mathrm{GU}(A)=\mathrm{GU}_{N}(q)$ when $\epsilon=-$, we see that the dimension of the $\lambda$-eigenspace is the same as that of the $\lambda^{\epsilon q}$-eigenspace. So if $\mathsf{supp}(g)<N/2$, then $g$ has a unique primary eigenvalue $\lambda$ which necessarily belongs to $\mu_{q-\epsilon}\subseteq\mathbb{F}_{q^{2}}^{\times}$. In this paper, we are primarily interested in the case $\epsilon=-$, but in §3, we consider both cases, with an eye to possible future applications.
We choose $x$ to be a real and regular semisimple element whose characteristic polynomial has a bounded number of irreducible factors. For example, in proving the theorem for $N=2p$, where $p$ is prime, as carried out in §4, we choose $x$ to have eigenvalues $\lambda^{\pm 1},\lambda^{\pm q},\ldots,\lambda^{\pm q^{p-1}}$, where $\lambda\in\mathbb{F}_{q^{p}}$ has degree $p$ over $\mathbb{F}_{q}$. This enables us to get strong upper bounds on $|\chi(x)|$ for all $\chi$ and good control over which $\chi$ satisfy $\chi(x)\neq 0$. Similarly, in §5, we prove the theorem for $N=2p+1$, using the same element $x$ as before but regarded as belonging to $\mathrm{SU}_{2p+1}(q)$ via the natural inclusion $\mathrm{SU}_{2p}(q)<\mathrm{SU}_{2p+1}(q)$.
In general, it is important to know which unipotent characters are non-zero on $x$, and this reduces to questions about characters on $\mathsf{S}_{N}$, which are discussed in §2. These results are used in §6 to prove the main theorem in the case that $N$ is of the form $2p+2r+\kappa$, where $p$ and $r$ are prime, $\kappa\in\{0,1\}$, and $r$ lies in a fixed interval as $N\to\infty$.
It is generally most difficult to prove that $g\in C^{2}$ when $\mathsf{supp}(g)$ is small, essentially because we can expect $|\chi(g)|$ to be small compared to $\chi(1)$ when the centralizer of $g$ is not too large. This is formalized in the uniform exponential character bounds of [7, 15]. These bounds are good enough that we need only consider $g$ with $\mathsf{supp}(g)$ less than an upper bound determined by the number of irreducible factors in the characteristic polynomial of $x$. The case of support 00 is of course particularly important, which is another way of saying that we must choose $x$ to be real. The case of support $1$ elements also needs special treatment; in fact, the treatment of transvections relies on a positivity property, see (4.8), which holds only for unitary groups, but not for linear groups.
For the remaining elements $g$, of support $\geq 2$, we break the sum (1.1) into two pieces, roughly speaking, a sum over the characters of low degree with $\chi(x)\neq 0$ and a sum over the characters of high degree with the same property. For the low degree characters $\chi$ (and when $g$ has not too large support), we have a very precise estimate for $\chi(g)/\chi(1)$, given in §3, which depends on Howe’s theory of dual reductive pairs [10] and may be of interest beyond its relevance to Thompson’s conjecture. For the high degree characters, $|\chi(g)|/\chi(1)$ is very small; in fact, for elements of support $\geq 2$, the centralizer bound is already good enough to handle the largest characters of $G$.
To conclude the proof, in §7 we employ a trick introduced in [15], to reduce from general $N$ to one of the cases treated in sections §4–§6, which involves choosing $x$ in block diagonal form. This is described in more detail at the beginning of the section.
To fully establish Thompson’s conjecture for all sufficiently large simple groups, it remains to treat orthogonal groups in odd characteristic, cf. [15, Theorem 7.7]. In principle, the approach of the current paper is also applicable to that case as well. This would require analogues for types $B$ and $D$ of several intermediate results in the paper, especially of Theorem 2, which still seem beyond our reach.
2. Some $\mathsf{S}_{N}$-character values
Let $N$ be a positive integer. For each partition $\lambda\vdash N$, let $\chi_{\lambda}$ denote the irreducible character of $\mathsf{S}_{N}$ associated to $\lambda$.
We identify a partition $\lambda\vdash N$ with its Ferrers diagram, with row lengths $\lambda_{1},\lambda_{2},\ldots$. Each box in the diagram has a row number $y\geq 1$ and a column number $x\geq 1$ and can be identified by the pair $(x,y)$. The layer number is the smaller of the two, and if the two are equal, we say the box is diagonal. By the $k^{\mathrm{th}}$ layer, we mean the set of boxes with layer number $k$; the size, that is, the number of boxes, of the $k^{\mathrm{th}}$ layer decreases monotonically with increasing $k$. The number of diagonal boxes equals the number of (non-empty) layers. The distance between the boxes $(x_{1},y_{1})$ and $(x_{2},y_{2})$ is $|x_{1}-x_{2}|+|y_{1}-y_{2}|$.
A border strip in $\lambda$ is a connected skew diagram containing no $2\times 2$ squares. By connectedness, the size of any border strip containing two boxes must be greater than the distance between the boxes. The maximum length of a border strip in $\lambda$ is $\lambda_{1}+\lambda^{\prime}_{1}-1$, where $\lambda^{\prime}$ is the transpose of $\lambda$. If the border strip does not include the last box in the first row of $\lambda$, its size can be no greater than $\lambda_{1}+\lambda^{\prime}_{1}-2$.
A border strip can be thought of as the union of its rows, which must be contained in a consecutive set of rows of $\lambda$, with the property that the first box of row $i$ lies in the same column as the last box of row $i+1$, or as the union of its columns, which must be contained in a consecutive set of columns of $\lambda$ such that the first box of column $i$ lies in the same row as the last box of column $i+1$.
A border strip tableau of type $(\alpha_{1},\ldots,\alpha_{m})$ of $\lambda$ is a labelling of its boxes by integers in $[1,m]$ such that the labels are non-decreasing in each row and column and such that the boxes labeled by any given $i$ form a border strip of size $\alpha_{i}$. Note that if $x_{1}<x_{2}$ and $y_{1}<y_{2}$ then $(x_{1},y_{1})$ and $(x_{2},y_{2})$ cannot have the same label $i$, since otherwise, the border strip of boxes labeled $i$ would contain the rectangle of which $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are opposite corners. Thus labels are strictly increasing along diagonals of type $(1,1)$. Regarding the last border strip in a border strip tableau of $\lambda$ as a union of columns, we see that if it contains the last box in the first row of $\lambda$, it is uniquely determined by $\alpha_{m}$.
The height of a tableau $T$ is the sum of the heights of its border strips, where the height of a single border strip is one less than the number of rows it meets. By the Murnaghan-Nakayama rule, the value of $\chi_{\lambda}$ at a cycle $x$ with orbit lengths $\alpha_{1},\ldots,\alpha_{m}$ is $\sum_{T}(-1)^{0pt(T)}$, where the sum is taken over border strip tableaux of type $(\alpha_{1},\cdots,\alpha_{m})$. The order of the $\alpha_{i}$ in this $m$-tuple may affect the number of terms in the sum but not, of course, the value of the sum.
We say that $\lambda$ is minimal for an element $x$ if there is some ordering of the orbit lengths of $x$ so that the number of tableaux of type $(\alpha_{1},\ldots,\alpha_{m})$ is at most $1$. This guarantees
$\chi_{\lambda}(x)\in\{-1,0,1\}.$
Proposition 2.1.
If $x\in\mathsf{S}_{N}$ has orbits of lengths $\alpha_{1}\leq\cdots\leq\alpha_{m}$ and
$\alpha_{i+1}\geq 2(\alpha_{1}+\cdots+\alpha_{i})-1$
for all $1\leq i<m$, then every $\lambda\vdash N$ is minimal for $x$. In particular, every irreducible character takes value 00 or ±$\pm 1$ on $x$.
Proof.
Since the transpose of any border strip tableau of $\lambda$ is a border strip tableau of the transpose partition $\lambda^{\prime}$, we may assume without loss of generality that the number $\lambda_{1}$ of columns is at least as great as the number $\lambda^{\prime}_{1}$ of rows.
If there are no boxes in the first row which are labeled $m$, then the last border strip is entirely contained in a rectangle with $\lambda_{1}$ columns and $\lambda^{\prime}_{1}-1$ rows, so
$\alpha_{m}\leq\lambda_{1}+\lambda^{\prime}_{1}-2\leq 2\lambda_{1}-2.$
Also,
$\alpha_{m}\leq\lambda_{2}+\lambda_{3}+\cdots.$
Adding twice this inequality to the previous inequality, we obtain
$23\alpha_{m}\leq 2(\lambda_{1}+\lambda_{2}+\cdots)-2=2N-$ 2 .
On the other hand,
$13\alpha_{m}=2\alpha_{m}+\alpha_{m}\geq 2\alpha_{m}+(2\alpha_{1}+\cdots+2\alpha_{m-1})-1=2N-$ 1 .
The contradiction implies that the first row must have some box labeled $m$, and this implies that the last box in the first row is so labeled. We conclude that the $m$th border strip is uniquely determined by $\alpha_{m}$, and this determines the partition obtained from $\lambda$ by removing the $m$th border strip. By induction on $m$, there can be at most one border strip tableau of $\lambda$ of type $(\alpha_{1},\ldots,\alpha_{m})$. ∎
Let $a<N/3$ be a positive integer, and let
[…]
Proposition 2.2.
Every partition $\lambda\vdash N$ is minimal for the element $x$ in (2.1). The total number of $\lambda$ such that $\chi_{\lambda}(x)\neq 0$ is $a(N-a)$. For any fixed $b$, the number of $\lambda$ for which $\lambda_{1}=N-b$ and $\chi_{\lambda}(x)\neq 0$ is at most $21+a(a+1)/$ 2. If $0\leq b<a$, the only $\lambda$ with $\lambda_{1}=N-b$ and $\chi_{\lambda}(x)\neq 0$ is the hook partition $\lambda=(N-b,1,\ldots,1)$.
Proof.
By Proposition 2.1, $\lambda$ is always minimal for $x$. Therefore,
$a(N-a)=|\mathbf{C}_{\mathsf{S}_{N}}(x)|=\sum_{\chi\in\mathrm{Irr}(\mathsf{S}_{N})}|\chi(x)|^{2}=|\{\chi\in\mathrm{Irr}(\mathsf{S}_{N})\mid\chi(x)\neq 0\}|.$
Next, we fix $b=N-\lambda_{1}$. Each border strip can contain at most one diagonal box, so $\chi_{\lambda}(x)=0$ unless $\lambda$ has at most two layers. Assuming there are $\leq 2$ layers, the size $N-k$ of the first layer is at least $N-a$, and, given $b$, this size determines the boxes in the first layer. For $k=0$, this determines $\lambda$ uniquely. Otherwise, there are at most $k$ possibilities for the second layer, so the total number of possibilities is at most $21+a(a+1)/$ 2.
Henceforth, we assume $b<a$. Suppose first that $\lambda$ has one layer, which implies that $\lambda=(N-b,1,\ldots,1)$. In any border tableau of type $(a,N-a)$, the unique diagonal box is labeled $1$, and the boxes labeled $2$ form a connected set, which implies that either they all belong to the first row or all to the first column. However, the first column has $b+1\leq a<N/3<N-a$ boxes, so the boxes labeled $2$ must be exactly the last $N-a$ boxes of the first row. Note that the first row has $N-b>N-a$ boxes, so this is, indeed, possible. Therefore, $\chi_{\lambda}(x)=\pm 1$.
If $\lambda$ has two layers, in any border strip tableau of type $(N-a,a)$, the second diagonal box must be labeled $2$. It is impossible for the last box in the first row to be labeled $2$ as well, since its distance from the second diagonal box is $N-b>a$. Therefore, all $N-b$ boxes in the first row are labeled $1$, which is possible only if $b\geq a$; otherwise, there are no tableaux of this shape at all, and $\chi_{\lambda}(x)=0$. ∎
Proposition 2.3.
Every $\lambda\vdash N$ is minimal for the element $y$ in (2.1). The number of $\lambda$ such that $\chi_{\lambda}(x)\neq 0$ is $(a-1)(N-a)$, unless $a=2$, in which case it is $2(N-2)$. For fixed $b$, the number of partitions $\lambda$ with $\lambda_{1}=N-b$ and $\chi_{\lambda}(y)\neq 0$ is at most $a^{2}$. If $10\leq b<a-$ 1, the only partition $\lambda$ with $\lambda_{1}=N-b$ and $\chi_{\lambda}(y)\neq 0$ is $(N)$ if $b=0$ and $(N-b,2,1,\ldots,1)$ if $b\geq 2$.
Proof.
By Proposition 2.1, $\lambda$ is always minimal for $y$. Therefore, assuming $a\geq 3$
$(a-1)(N-a)=|C_{\mathsf{S}_{N}}(y)|=\sum_{\chi\in\mathrm{Irr}(\mathsf{S}_{N})}|\chi(y)|^{2}=|\{\chi\in\mathrm{Irr}(\mathsf{S}_{N})\mid\chi(y)\neq 0\}|.$
If $a=2$, then $|C_{\mathsf{S}_{N}}(y)|=2(N-2)$.
Next, we assume that $b$ is fixed, and $\lambda_{1}=N-b$. Let $d$ be the number of layers of $\lambda$. If $d>3$, there are no border strip tableaux consisting of three strips, so we may assume $d\leq 3$. As in the proof of Proposition 2.2, we fix $k\in[0,a]$ to be the size of the complement of the first layer of $\lambda$, so $k$ and $b$ together determine the first layer. If $d\leq 2$, there are $\leq k$ possibilities for the second layer. If $d=3$ and $\lambda$ has a tableau of type $(a-1,N-a,1)$, then the third diagonal box must be labeled $3$, so all boxes in the third layer must be so labeled, meaning that the third layer must consist only of its diagonal box. There are $\leq k-1$ possibilities for the second layer in this case. In total, therefore, for any given $k$, there are at most $12k-$ 1 possibilities for $\lambda$ with $\chi_{\lambda}(y)\neq 0$; summing over $k\in[0,a]$, there are $\leq a^{2}$ possibilities for $\lambda$.
Finally, we assume that $b<a-1$. The result is trivial for $b=0$, so we assume $b>0$.
Case $d=1$.
As $b<a-1$, there is a unique border strip of size $N-a$, namely the last $N-a$ entries in the first row. The complement of this border strip is the Ferrers diagram of $(a-b,\underbrace{1,\ldots,1}_{b})$. This cannot contain a border strip of size $a-1$ because we have already assumed that $b>0$ and $a-b>1$. Therefore, $\lambda$ is minimal for $y$, and, in fact, $\chi_{\lambda}(y)=0$.
Case $d=2$.
We claim that the only way in which a border tableau of type $(1,a-1,N-a)$ can exist is if the second layer has size $1$. Note that $b$ is the number of boxes which do not belong to the first row. Therefore, the condition $b<a-1$ implies that the number of boxes in the first row and labeled by $2$ is greater than the number of boxes not in the first row and labeled by $3$. Given any box $(1,y)$ labeled by $2$, the box $(2,y+1)$, if it exists, must be labeled by $3$. Therefore, there exists some box $(1,y)$ labeled by $2$ such that $\lambda_{2}\leq y$. This is only possible if the first box in the first row which is labeled $3$ is the only box in its column in $\lambda$, which, by connectedness, means that the third border strip consists of the last $N-a$ boxes in the first row. The complement of this border strip has a tableau of type $(1,a-1)$. Since the longest possible border strip in any Ferrers diagram has length equal to the size of the first layer, this is possible only if the second layer consists only of the $(2,2)$ box.
Case $d=3$.
Let $T$ be any tableau of type $(N-a,a-1,1)$. The existence of $T$ implies there is only one box in the third layer. Removing this box, we obtain a partition $\mu$ with two layers, and $T$ gives a border strip tableau of $\mu$ of type $(N-a,a-1)$. As $b<a-1$, by the last paragraph of the proof of Proposition 2.2, there is no such tableau. ∎
3. A sharp asymptotic for character ratios in groups of type $A$
Fix ϵ=±italic-$\epsilon=\pm$, $q$, and $N$, and let $G:=G_{N}:=\mathrm{GL}_{N}^{\epsilon}(q)$. Recall the notion of the true level $\mathfrak{l}^{*}(\chi)$ of any $\chi\in\mathrm{Irr}(G)$ introduced in [7, Definition 1]. Let $\tau_{N}$ denote the reducible Weil character of $G$:
$\tau_{N}(g)=\epsilon^{N}(\epsilon q)^{\dim\ker(g-1)},$
(in which the sign $\epsilon$ is interpreted as an element of {±$\{\pm 1\}$). Then $\mathfrak{l}^{*}(\chi)$ is the smallest integer $j\geq 0$ such that $\chi$ is an irreducible constituent of $\tau^{j}$.
Our first main result visibly improves on [14, Proposition 10.5]:
Theorem 2.
Let $m$ and $j$ be fixed positive integers, and let ϵ=±italic-$\epsilon=\pm$. If
$N\geq 3j^{2}+j+2m(j+1)+5,$
$\chi$ is an irreducible complex character of $G=\mathrm{GL}^{\epsilon}_{N}(q)$ of true level $j$, and $g\in G$ has support $m$ and primary eigenvalue $\lambda$, then
(3.1) $\Bigm{|}\frac{(\epsilon q)^{mj}\chi(g)}{\chi(1)}\cdot\frac{\overline{\chi(\lambda\cdot 1_{V})}}{\chi(1)}-1\Bigm{|}<q^{-N/2}.$
Proof.
We will follow in part the proof of [7, Theorem 1.6]. Let $j^{2}+j+1\leq N$, and let $\chi\in\mathrm{Irr}(G)$ have $\mathfrak{l}^{*}(\chi)=j\geq 1$. Let $S:=G_{j}$, and let $\tau:=\tau_{jN}$. The restriction of $\tau$ from $G_{jN}$ to $G_{N}\times G_{j}$ gives a decomposition
$\tau|_{G\times S}=\sum_{\alpha\in\mathrm{Irr}(S)}D_{\alpha}\boxtimes\alpha,$
where $D_{\alpha}$ is given by the well-known formula
(3.2) $D_{\alpha}(g)=\frac{1}{|S|}\sum_{s\in S}\tau(g\otimes s)\bar{\alpha}(s).$
By [7, Theorem 1.1], there exists a unique irreducible factor $D_{\alpha}^{\circ}$ which is of true level $j$, and all other factors have strictly lower true level. Moreover, for each $\chi$ of true level $j$, there exists a unique $\alpha\in\mathrm{Irr}(S)$ such that $\chi=D_{\alpha}^{\circ}$. Consider any $g\in G\smallsetminus\mathbf{Z}(G)$ of support $m$ and primary eigenvalue $\lambda$.
If $\epsilon=+$ (resp. $\epsilon=-$), we define $Q:=q$ (resp. $Q:=q^{2}$). We write $A:=\mathbb{F}_{Q}^{N}$, $B:=\mathbb{F}_{Q}^{j}$, $V:=A\otimes_{\mathbb{F}_{Q}}B$, which is endowed with the tensor product of the Hermitian forms on $A$ and $B$ if $\epsilon=-$. We will bound $D^{\circ}_{\alpha}(1)$ and $|D^{\circ}_{\alpha}(g)|$ using (3.2). According to Definition 3.2 and Theorem 1.1 of [7], embedding $G=G\times\{1\}\hookrightarrow G\times S\hookrightarrow G_{jN}$, we can write
(3.3) $\tau|_{G}=\sum^{M}_{i=1}a_{i}\theta_{i},~{}~{}D^{\prime}_{\alpha}:=D_{\alpha}-D^{\circ}_{\alpha}=\sum^{M^{\prime}}_{i=1}b_{i}\theta_{i},$
where $\theta_{i}\in\mathrm{Irr}(G)$ are pairwise distinct, $a_{i},b_{i}\in\mathbb{Z}_{\geq 0}$, $M\geq M^{\prime}$, $a_{i}\geq b_{i}$ if $i\leq M^{\prime}$, $\mathfrak{l}^{*}(\theta_{i})\leq j$ for all $i$. In fact, if $i\leq M^{\prime}$, then $\mathfrak{l}^{*}(\theta_{i})\leq j-1$, and so $\mathfrak{l}(\theta_{i})=\mathfrak{l}^{*}(\theta_{i})\leq j-1<N/2$ as $j\leq N/2$, whence $\theta_{i}(1)\leq q^{N(j-1)}$ by [7, Theorem 1.2].
Let $k(X)=|\mathrm{Irr}(X)|$ denote the class number of a finite group $X$. By [4, Proposition 3.5], $k(\mathrm{GL}_{n}(q))\leq q^{n}$. Note that $M^{\prime}$ cannot exceed the total number of irreducible characters of true level $<j$. Hence, when $\epsilon=+$ by [7, Theorem 1.1] we have
$M^{\prime}\leq\sum^{j-1}_{i=0}k(\mathrm{GL}_{i}(q))\leq\sum^{j-1}_{i=0}q^{i}<q^{j}.$
Also, $\sum^{M}_{i=1}a_{i}^{2}=[\tau|_{G},\tau|_{G}]_{G}=[\tau^{2}|_{G},1_{G}]_{G}\leq 8q^{j^{2}}$ by [7, Lemma 2.4] (since $\tau$ is the permutation character of $\mathrm{GL}_{jN}(q)$ on $\mathbb{F}_{q}^{Nj}$, and therefore $\tau^{2}|_{G}$ is the permutation character of $G$ acting on $\mathbb{F}_{q}^{Nj}\oplus\mathbb{F}_{q}^{Nj}$). It follows that
$(\sum^{M^{\prime}}_{i=1}b_{i})^{2}\leq M^{\prime}\sum^{M^{\prime}}_{i=1}b_{i}^{2}\leq q^{j}\sum^{M}_{i=1}a_{i}^{2}\leq 8q^{j^{2}+j},$
and so
(3.4) $D^{\prime}_{\alpha}(1)\leq\sum^{M^{\prime}}_{i=1}b_{i}q^{N(j-1)}\leq\sqrt{8q^{j^{2}+j}}q^{N(j-1)}\leq q^{N(j-1)+(j^{2}+j+3)/2}.$
Assume now that $\epsilon=-$. Then every irreducible constituent of $D^{\prime}_{\alpha}$ has level $\leq j-2$ by [7, Corollary 4.8], and $k(\mathrm{GU}_{n}(q))\leq 8.26q^{n}$ by [4, §3.3]. So in this case we have
$M^{\prime}\leq\sum^{j-2}_{i=0}k(\mathrm{GU}_{i}(q))\leq 8.26\sum^{j-2}_{i=0}q^{i}<8.26q^{j-1}.$
Also, $\sum^{M}_{i=1}a_{i}^{2}=[\tau|_{G},\tau|_{G}]_{G}=[\tau^{2}|_{G},1_{G}]\leq 2q^{j^{2}}$ by [7, Lemma 2.4] (since $\tau^{2}$ is the permutation character of $\mathrm{GU}_{jN}(q)$ on $\mathbb{F}_{q^{2}}^{Nj}$). It follows that
$(\sum^{M^{\prime}}_{i=1}b_{i})^{2}\leq M^{\prime}\sum^{M^{\prime}}_{i=1}b_{i}^{2}\leq 8.26q^{j-1}\sum^{M}_{i=1}a_{i}^{2}\leq 16.52q^{j^{2}+j-1}\leq 8.26q^{j^{2}+j},$
and so
$D^{\prime}_{\alpha}(1)\leq\sum^{M^{\prime}}_{i=1}b_{i}q^{N(j-2)}\leq\sqrt{8.26q^{j^{2}+j}}q^{N(j-2)}<\sqrt{8q^{j^{2}+j}}q^{N(j-1)},$
and thus (3.4) holds in this case as well.
Recall that $g$ has support $m<N/2$, which implies that $g$ has a primary eigenvalue $\lambda$ with $\lambda^{q-\epsilon}=1$, and the dimension $\delta_{A}(g)$ of the largest eigenspace of $g$ on $A$, which corresponds to $\lambda$, satisfies
$\delta_{A}(g)=N-m>N/2.$
Multiplying $g$ by the central element $\lambda^{-1}I_{A}$, we may assume $\lambda=1$. Consider any $s\in S=\mathrm{GL}^{\epsilon}(B)\cong\mathrm{GL}^{\epsilon}_{j}(q)$. By [7, Lemma 8.2] we have
$d_{V}(g\otimes s):=\dim\mathrm{Ker}(g\otimes s-1_{V})\leq j(N-m).$
In fact, this bound is attained for $s_{0}=1_{B}$. For all other elements $s\neq s_{0}$ in $S$ we have the stronger bound
(3.5) $d_{V}(g\otimes s)\leq(N-m)(j-2)+N\leq N^{*}:=\max(m,N(j-1)).$
Note that $\tau(g\otimes s)=\epsilon^{Nj}(\epsilon q)^{d_{V}(g\otimes s)}$; in particular,
$\tau(g\otimes s_{0})=\epsilon^{Nj}(\epsilon q)^{(N-m)j}=\epsilon^{mj}q^{(N-m)j}.$
By (3.2) we can now write
(3.6) $\chi(g)=D^{\circ}_{\alpha}(g)=\frac{\alpha(1)}{|S|}\Bigl{(}\epsilon^{mj}q^{(N-m)j}+X\Bigr{)},$
where
$X:=\sum_{s_{0}\neq s\in S}\frac{\tau(g\otimes s)\bar{\alpha}(s)}{\alpha(1)}-\frac{|S|}{\alpha(1)}D^{\prime}_{\alpha}(g).$
Since $D^{\prime}_{\alpha}$ is either zero or a true character, we have $|D^{\prime}_{\alpha}(g)|\leq D^{\prime}_{\alpha}(1)$. Furthermore,
$|S|=|\mathrm{GL}^{\epsilon}_{j}(q)|\leq(q+1)q^{j^{2}-1}<q^{j^{2}+0.6}.$
Using (3.4) and (3.5), we have
$|X|\leq Z:=q^{N^{*}+j^{2}+0.6}+q^{N(j-1)+(3j^{2}+j)/2+2.1}.$
Similarly, $d_{V}(I_{A}\otimes s)=Nd_{V}(s)\leq N(j-1)$ for $s\neq s_{0}$, and so
(3.7) $\chi(1)=D^{\circ}_{\alpha}(1)=\frac{\alpha(1)}{|S|}\Bigl{(}q^{Nj}+Y\Bigr{)},$
where
$Y:=\sum_{s_{0}\neq s\in S}\frac{\tau(I_{A}\otimes s)\bar{\alpha}(s)}{\alpha(1)}-\frac{|S|}{\alpha(1)}D^{\prime}_{\alpha}(1)$
and $|Y|\leq|Z|$ as well. Now, (3.6) and (3.7) imply
$\Bigm{|}\frac{(\epsilon q)^{mj}\chi(g)}{\chi(1)}-1\Bigm{|}=\Bigm{|}\frac{q^{Nj}+X(\epsilon q)^{mj}}{q^{Nj}+Y}-1\Bigm{|}=\Bigm{|}\frac{X(\epsilon q)^{mj}-Y}{q^{Nj}+Y}\Bigm{|}\leq\frac{Z(q^{mj}+1)}{q^{Nj}-Z}.$
Hence to prove (3.1), it suffices to prove
$Z(q^{mj+N/2}+q^{N/2}+1)<q^{Nj}.$
As $mj\geq 1$ and $N/2\geq 6$, to prove the latter inequality it suffices to show
(3.8) $Zq^{mj+N/2+0.6}\leq q^{Nj}.$
Suppose $j\geq 2$. Then $N^{*}=N(j-1)$ in (3.5); furthermore,
$q^{(j^{2}+j+3)/2}+1<q^{(j^{2}+j)/2+1.8}.$
Hence
$Z=q^{N(j-1)+(3j^{2}+j)/2+2.1}+q^{N^{*}+j^{2}+0.6}<q^{N(j-1)+(3j^{2}+j)/2+2.4}$
and so (3.8) holds when $N\geq 3j^{2}+j+2mj+6$. On the other hand, if $j=1$ and $N\geq 4m+9$, then $N^{*}=m$, $Z=q^{m+1.6}+q^{4.1}$ and so (3.8) holds as well. ∎
We will also need the following quantitative statement, which is well known in the $\mathrm{GL}$-case:
Lemma 3.1.
Let $q$ be any prime power, $N\in\mathbb{Z}_{\geq 6}$, and $g$ any element in $G=\mathrm{GU}_{N}(q)$. If $g$ has support $1$, then
$|\mathbf{C}_{G}(g)|\leq q^{N^{2}-2N+3.2}.$
If $g$ has support $\geq 2$, then
$|\mathbf{C}_{G}(g)|\leq q^{N^{2}-3.4N+5}.$
Proof.
We will use the estimates
(3.9) $q^{N^{2}}<|\mathrm{GU}_{N}(q)|<q^{N^{2}+0.6}$
for any $N\geq 1$. Let $m:=\mathsf{supp}(g)$, and consider the case $m\geq 4$. By [15, Proposition 4.2(a)], $\mathbf{C}_{\mathrm{GL}_{N}(\bar{\mathbb{F}}_{q})}(g)$ has dimension $\leq N(N-m)$. Now we can follow the proof of [15, Proposition 6.1] in passing to the finite case and increase the upper bound by a factor of $q^{0.6N}$ to account for the upper bound in (3.9). This shows
$|\mathbf{C}_{G}(g)|\leq q^{N(N-m+0.6)}\leq q^{N^{2}-3.4N}.$
In the remaining cases $1\leq m\leq 3$, let $s$ denote the semisimple part of $g$. Suppose first that $s\in\mathbf{Z}(G)$. Then we may assume that $g$ is unipotent, and use [16, Theorem 7.1] and (3.9) to bound $|\mathbf{C}_{G}(g)|$. Denoting the Jordan block of size $i$ and eigenvalue $1$ by $J_{i}$, we have the following possibilities for $g$. If $g=J_{1}^{N-5}\oplus J_{2}^{1}\oplus J_{3}^{1}$, then
$|\mathbf{C}_{G}(g)|\leq q^{N^{2}-6N+14+1.8}\leq q^{N^{2}-3.4N+5}$
when $N\geq 5$ (and the term $1.8$ accounts for three $\mathrm{GU}$-factors in $\mathbf{C}_{G}(g)$).
Similarly, if $g=J_{1}^{N-4}\oplus J_{4}^{1}$, then $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-6N+13.2}$.
If $g=J_{1}^{N-3}\oplus J_{3}^{1}$, then $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-4N+7.2}$.
If $g=J_{1}^{N-6}\oplus J_{2}^{3}$, then $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-6N+19.2}<q^{N^{2}-3.4N+5}$ since $N\geq 6$.
If $g=J_{1}^{N-4}\oplus J_{2}^{2}$, then $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-4N+9.2}\leq q^{N^{2}-3.4N+5}$ since $N\geq 7$.
If $g=J_{1}^{N-2}\oplus J_{2}^{1}$, then $m=1$, and $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-2N+3.2}$.
Now we may assume $s\notin\mathbf{Z}(G)$. Then $1\leq\mathsf{supp}(s)\leq m\leq 3$. If $\mathsf{supp}(s)=3$, then as $N\geq 7$ we have
$|\mathbf{C}_{G}(g)|\leq|\mathbf{C}_{G}(s)|\leq|\mathrm{GU}_{3}(q)\times\mathrm{GU}_{N-3}(q)|\leq q^{N^{2}-6N+19.2}<q^{N^{2}-3.4N+5}.$
Similarly, if $\mathsf{supp}(s)=2$, then
$|\mathbf{C}_{G}(g)|\leq|\mathbf{C}_{G}(s)|\leq|\mathrm{GU}_{2}(q)\times\mathrm{GU}_{N-2}(q)|\leq q^{N^{2}-4N+9.2}\leq q^{N^{2}-3.4N+5}.$
In the remaining cases, $\mathsf{supp}(s)=1$ and $\mathbf{C}_{G}(s)=\mathrm{GU}_{1}(q)\times\mathrm{GU}_{N-1}(q)$. If $m=1$, then $|\mathbf{C}_{G}(g)|=|\mathbf{C}_{G}(s)|\leq q^{N^{2}-2N+3.2}$. If $2\leq m\leq 3$, then the unipotent part of the $\mathrm{GU}_{N_{1}}(q)$-component of $g$ has support $1$ or $2$, respectively, and $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-4N+7.8}$, respectively $|\mathbf{C}_{G}(g)|\leq q^{N^{2}-6N+15.8}$. ∎
4. Thompson’s conjecture for $\mathrm{PSU}_{2p}(q)$
Let $p\geq 3$ be any prime, and let $q\leq 7$ be a prime power (which may be coprime to $p$). In this section we show that Thompson’s conjecture holds for the simple group $\mathrm{PSU}_{2p}(q)$ as long as $p$ is sufficiently large. Together with the main result of [3], this implies Thompson’s conjecture for all sufficiently large unitary simple groups in the case that the rank is twice a prime less $1$.
We will work inside the general unitary group $G=\mathrm{GU}_{2p}(q)$ and consider any element $x\in G$ which has some fixed $\xi\in\mathbb{F}_{q^{p}}\smallsetminus\mathbb{F}_{q}$ as an eigenvalue. Note that $\xi$ has degree $p$ over $\mathbb{F}_{q}$, and the orbit of $\xi$ under the map $\lambda\mapsto\lambda^{-q}$ has length $2p$ and contains $\xi^{-1}$. It follows that $x$ is regular semisimple and real and belongs to $\mathrm{SU}_{2p}(q)$.
Theorem 3.
There is an explicit absolute constant $A>0$ such that the following statement holds when the prime $p$ is at least $A$. If $C$ denotes the conjugacy class of the image of the element $x$ described above in the simple group $S=\mathrm{PSU}_{2p}(q)$, then $S=C^{2}$. In fact, $(x^{\mathrm{SU}_{2p}(q)})^{2}$ contains all unipotent elements and all elements of support $\geq 2$ in $\mathrm{SU}_{2p}(q)$.
The rest of the section is devoted to the proof of Theorem 3. The chosen element $x$ has
$T:=\mathbf{C}_{G}(x)\cong C_{q^{2p}-1}.$
In particular, $T\cap\mathrm{SU}_{2p}(q)$ has index $q+1$ in $T$, and hence the conjugacy class $x^{G}$ is a single $\mathrm{SU}_{2p}(q)$-conjugacy class. As $x$ is real, it suffices to show that, any non-central element $g\in\mathrm{SU}_{2p}(q)$ is a product of two $G$-conjugates of $x$, except possibly in the case $g=zh$ where $1\neq z\in\mathbf{Z}(\mathrm{SU}_{2p}(q))$ and $h$ is a transvection. By the extension [9, Lemma 5.1] of Gow’s lemma [6], the statement holds in the case $g$ is semisimple. So we may assume that $g$ is not semisimple, and furthermore $g$ is a transvection if $\mathsf{supp}(g)=1$.
As $x$ is real, we can rewrite (1.1) as follows:
(4.1) $\sum_{\chi\in\mathrm{Irr}(G)}\frac{|\chi(x)|^{2}\bar{\chi}(g)}{\chi(1)}\neq 0.$
The summation in (4.1) includes $q+1$ linear characters which all take value $1$ at any element in $\mathrm{SU}_{2p}(q)$. Hence it suffices to show
(4.2) |*$\Bigm{|}\sum\nolimits^{*}\frac{|\chi(x)|^{2}\chi(g)}{\chi(1)}\Bigm{|}<q+1,$
where $\sum^{*}$ indicates the sum over $\chi\in\mathrm{Irr}(G)$ such that $\chi(1)>1$, $\chi(x)\neq 0$, and $\chi(g)\not\in[0,\infty)$. We denote by $\Sigma^{*}$ the sum in (4.2).
We will identify the dual group $G^{*}$ with $G$. Note that any non-central semisimple element $s\in G$ has the property that $\mathbf{C}_{G}(s)$ is a Levi subgroup $L$ of $G$. Fix an embedding of $\bar{\mathbb{F}}_{q}^{\times}$ into $\mathbb{C}^{\times}$. Then one can identify $\mathbf{Z}(\mathbf{C}_{G}(s))$ with
$\mathrm{Hom}(\mathbf{C}_{G}(s)/[\mathbf{C}_{G}(s),\mathbf{C}_{G}(s)],\mathbb{C}^{\times})$
as in [5, (1.16)], and the linear character of $L=\mathbf{C}_{G}(s)$ corresponding to $s$ will be denoted by $\hat{s}$. Now, any irreducible character $\chi$ in the Lusztig series labeled by $s$ is
(4.3) $\chi=\pm R^{G}_{L}(\hat{s}\psi),$
where $\psi$ is a unipotent character of $L$; see [5, p. 116]. Let $\mathsf{St}_{G}$ and $\mathsf{St}_{L}$ denote the Steinberg character of $G$, respectively of $L$. By [2, Corollary 10.2.10(ii)],
(4.4) $\mathsf{St}_{G}\cdot\chi=\pm\mathrm{Ind}^{G}_{L}(\hat{s}\psi\cdot\mathsf{St}_{L}).$
Assume now that $\chi(x)\neq 0$. As $x$ is regular semisimple,
$\mathsf{St}_{G}(x)=\pm 1.$
Hence (4.4) implies that, up to conjugacy, $x\in L=\mathbf{C}_{G}(s)$. It follows that $s\in\mathbf{C}_{G}(x)=T$; in fact, $s\in T\smallsetminus\mathbf{Z}(G)$. Now using the primality of $p$, we see that the characters $\chi\in\mathrm{Irr}(G)$ that are non-vanishing at $x$ are divided into the following types:
  1. (I)
    Unipotent characters, multiplied possibly by a linear character of $G$.
  2. (II)
    Characters $\chi$ in (4.3) with $s$ belonging to the unique subgroup $C_{q^{2}-1}$ of $T$, but not in $\mathbf{Z}(G)\cong C_{q+1}$ (which is possible only if $q>2$). For any such $s$, $\mathbf{C}_{G}(s)\cong\mathrm{GL}_{p}(q^{2})$, a rational Levi subgroup of $G$.
  3. (III)
    Characters $\chi$ in (4.3) with $s$ belonging to the unique subgroup $C_{q^{p}+1}$ of $T$, but not in $C_{q^{2}-1}$. For any such $s$, $\mathbf{C}_{G}(s)\cong\mathrm{GU}_{2}(q^{p})$.
  4. (IV)
    Characters $\chi$ in (4.3) with $s$ not belonging to the union of the unique subgroups $C_{q^{p}+1}$ and $C_{q^{2}-1}$ of $T$. Any such $s$ is regular semisimple, with $\mathbf{C}_{G}(s)=T$.
Let $m$ denote the support of $g$, which is also $\mathsf{supp}(zg)$ for any $z\in\mathbf{Z}(G)$. We will use the character bound [15, Theorem 5.5]
(4.5) $\frac{|\chi(g)|}{\chi(1)}\leq\chi(1)^{-\sigma m/N}$
for an explicit absolute constant $\sigma>0$, and
$N:=2p.$
Lemma 4.1.
There is an absolute constant $N_{1}>0$ such that when $N\geq N_{1}$, the total contribution of characters of type (I) to $\Sigma^{*}$ has absolute value $<(q+1)/2$.
Proof.
It suffices to prove that the contribution of unipotent characters $\chi$ of type (I) to $\Sigma^{*}$ has absolute value $<1/2$. Since $x$ is regular semisimple with centralizer $T\cong C_{q^{N}-1}$, we can have $\chi(x)\neq 0$ only when the unipotent character $\chi$ is labeled by one of the $N$ hook partitions $\alpha=(N-j,1^{j})$, in which case
$|\chi(x)|=1,$
cf. the proof of [13, Proposition 3.1.5].
First we consider $\chi=\chi_{j}=\chi^{(N-j,1^{j})}$ with $j\geq\sqrt{N/13}$. Choosing $N_{1}$ sufficiently large, when $N\geq N_{1}$ we have $\mathfrak{l}^{*}(\chi)=j$ by [7, Theorem 3.9], and so, by [7, Theorem 1.2], $\chi(1)\geq q^{c_{1}N^{3/2}}$ for some absolute constant $c_{1}>0$. So, by (4.5), the total contribution of these unipotent characters to $|\Sigma^{*}|$ is at most
$Nq^{-c_{1}\sigma\sqrt{N}}<1/12$
when $N_{1}$ is large enough.
Now we look at the $\chi=\chi_{j}$ with $1\leq j\leq\sqrt{N/13}$. Suppose that
$m=\mathsf{supp}(g)\geq\sqrt{N/13}.$
Since $\chi(1)\geq q^{N/2}$, by (4.5), the total contribution of these unipotent characters to $|\Sigma^{*}|$ is at most
$\sqrt{N/13}q^{-\sigma\sqrt{N/52}}<1/3$
when $N_{1}$ is large enough, yielding the claim in this case.
In the remaining case, we have $m\leq\sqrt{N/13}$ and $j\leq\sqrt{N/13}$. By Theorem 2,
$\frac{\chi_{j}(g)}{\chi_{j}(1)}=\epsilon_{j}+(-q)^{-mj}\gamma_{\lambda},$
where $|\epsilon_{j}|<q^{-N/2}$ and $\gamma_{\lambda}=\chi(\lambda I_{A})/\chi(1)$ for the primary eigenvalue $\lambda$ of $g$. If $m\geq 2$, then
$\Bigm{|}\sum_{j\geq 1}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}$
is at most $\sum_{j\geq 1}4^{-j}=1/3$. If $m=1$, then $g$ is a transvection by our convention, so $\lambda=1$, $\gamma_{\lambda}=\gamma_{1}=1$, and
$\Bigm{|}\sum_{j\geq 1}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq 1/(q+1)\leq 1/3.$
Choosing $N_{1}$ large enough, we therefore have
$\Bigm{|}\sum_{1\leq j\leq\sqrt{N/13}}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq\frac{1}{3}+\frac{1}{24}$
and
$\Bigm{|}\sum_{1\leq j\leq\sqrt{N/13}}\epsilon_{j}\Bigm{|}\leq 2^{-N/2}\sqrt{N/13}\leq\frac{1}{24}.$
It follows that
$\Bigm{|}\sum_{1\leq j\leq\sqrt{N/13}}\frac{|\chi_{j}(x)|^{2}\chi_{j}(g)}{\chi_{j}(1)}\Bigm{|}\leq\frac{1}{3}+\frac{1}{12},$
and so we are done in this case as well. ∎
Lemma 4.2.
There is an absolute constant $N_{2}>0$ such that when $N\geq N_{2}$, the total contribution of characters of type (II) to $\Sigma^{*}$ has absolute value $<1/4$.
Proof.
Each of these characters can be written in the form (4.3), where $L=\mathrm{GL}_{p}(q^{2})$ is unique up to conjugacy in $G$. For the element $x$, we can use the bound
(4.6) $|\chi(x)|\leq N$
for all irreducible characters of $G$, which follows from [14, Corollary 7.6]. There are fewer than $q^{2}\leq 49$ choices for the element $s$ (up to conjugacy). Furthermore, by (4.4), the condition $\chi(x)\neq 0$ implies that (after conjugating $x$ suitably) $x\in L$ and $\psi(x)\neq 0$. Viewed as an element in $L$, $x$ is again regular semisimple with centralizer $C_{q^{2p}-1}$. Hence the unipotent character $\psi$ is labeled by one of the $N/2$ hook partitions of $p=N/2$. Thus there are fewer than $25N$ characters of type (II) with $\chi(x)\neq 0$. Each of them has degree
$\chi(1)\geq[G:L]_{q^{\prime}}=(q+1)(q^{3}+1)\ldots(q^{N-1}+1)>q^{N^{2}/4}.$
Hence, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$425N^{3}q^{-\sigma N/4}<1/$ 4
when $N\geq N_{2}$ and $N_{2}$ is large enough. ∎
To complete the proof of (4.2) and of Theorem 3, it remains to prove:
Proposition 4.3.
There is an absolute constant $N_{3}>0$ such that when $N\geq N_{3}$, the total contribution of characters of types (III) and (IV) to $\Sigma^{*}$ has absolute value $<1/4$.
Proof.
(a) First we consider the case $\mathsf{supp}(g)\geq 2$. By Lemma 3.1 we have
$|\chi(g)|\leq\sqrt{|\mathbf{C}_{G}(g)|}\leq q^{N^{2}/2-1.7N+2.5}$
for any $\chi$; and we still have the bound (4.6) for $|\chi(x)|$.
Each $\chi$ of type (III) can be written in the form (4.3), where $L=\mathrm{GU}_{2}(q^{p})$ is unique up to conjugacy in $G$, and
$\chi(1)\geq[G:L]_{q^{\prime}}=(q+1)(q^{2}-1)\ldots(q^{p-1}-1)\cdot(q^{p+1}-1)(q^{p+2}+1)\ldots(q^{2p-1}+1)>q^{N^{2}/2-N-1}$
by [11, Lemma 4.1(iii)]. There are fewer than $q^{N/2}$ choices for $s$, and two choices for the unipotent character $\psi$. Thus there are fewer than $2q^{N/2}$ choices for $\chi$ of type (III). Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$2N^{2}q^{N/2}q^{-0.7N+3.5}\leq N^{2}2^{3.5-0.2N}<\frac{1}{8}$
when $N\geq N_{3}$ and $N_{3}$ is large enough.
Each $\chi$ of type (IV) can be written in the form (4.3), where $L=\mathrm{GL}_{1}(q^{2p})$ is unique up to conjugacy in $G$, and
$\chi(1)\geq[G:L]_{q^{\prime}}=(q+1)(q^{2}-1)\ldots(q^{2p-1}+1)>q^{(N^{2}-N)/2}.$
There are fewer than $q^{N}$ choices for $s$, and one choice for the unipotent character $\psi$. Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$N^{2}q^{N}q^{-1.2N+2.5}\leq N^{2}2^{2.5-0.2N}<\frac{1}{8}$
when $N\geq N_{3}$ and $N_{3}$ is large enough.
(b) Next assume that $\mathsf{supp}(g)=1$. Even though (4.1) is already proved in the case $g$ is semisimple, for further use we will consider this case here. So assume that $g$ is semisimple with $\mathsf{supp}(g)=1$. Then
(4.7) $\mathbf{C}_{G}(g)=M\cong\mathrm{GU}_{1}(q)\times\mathrm{GU}_{N-1}(q).$
Consider a Levi subgroup $L=\mathrm{GU}_{2}(q^{p})$ and assume that $M\cap L$ contains a maximal torus $S$ of $G$. Then $S$ contains the central torus $\mathbf{Z}(L)\cong C_{q^{p}+1}$ of $L$. In particular, $S$ contains a generator $t$ of $\mathbf{Z}(L)$, all of whose eigenvalues on $\mathbb{F}_{q^{2}}^{N}$ are of order $q^{p}+1$. But then $t$ cannot belong to $M$, since any element in $M$ (namely its $\mathrm{GU}_{1}(q)$-component) has an eigenvalue of order dividing $q+1$, a contradiction.
The same argument shows that $M\cap L$ cannot contain any maximal torus of $L=\mathrm{GL}_{1}(q^{2p})$. By the Mackey formula, see [2, Theorem 5.2.1 and p. 140], we now have
${}^{*}\!R^{G}_{M}(\chi)=0$
for any character $\chi$ of type (III) or (IV). Together with (4.7), this implies by [21, Lemma 13.3] that $\chi(g)=0$ for all these characters.
(c) In the rest of the proof, we will assume that $g$ is a transvection, and aim to show that
(4.8) $\chi(g)=\frac{\chi(1)}{q^{N-1}+1}$
for any character $\chi$ of type (III) or (IV). In particular, $\chi(g)>0$, so it does not contribute to $\Sigma^{*}$ in (4.2).
To do this, we consider $G^{+}:=\mathrm{GL}_{2p}(q)$ (and identify its dual with $G^{+}$), and its Levi subgroups
$L_{1}^{+}:=L^{+}\cong\mathrm{GL}_{2}(q^{p}),~{}L_{2}^{+}:=T^{+}\cong\mathrm{GL}_{1}(q^{2p}),~{}M^{+}\cong\mathrm{GL}_{1}(q)\times\mathrm{GL}_{N-1}(q).$
Furthermore, we consider any character $\chi^{+}\in\mathrm{Irr}(G^{+})$ that lies in the Lusztig series labeled by a semisimple element $s^{+}$ with $\mathbf{C}_{G^{+}}(s^{+})=L_{i}^{+}$, $i=1,2$; in particular, $\chi^{+}$ is Lusztig induced from $L_{2}^{+}$ by (4.3). The same argument as in (b) shows that $L_{i}^{+}\cap M^{+}$ contains no maximal torus of $G^{+}$, and hence
(4.9) ${}^{*}\!R^{G^{+}}_{M^{+}}(\chi^{+})=0$
by the Mackey formula.
The Levi subgroup $M^{+}$ is a Levi subgroup of a parabolic subgroup $P=U\rtimes M^{+}$ of $G^{+}$ (the stabilizer of a line in $\mathbb{F}_{q}^{N}$), where the unipotent radical $U$ is an elementary abelian subgroup of order $q^{N-1}$. Furthermore, $M^{+}$ acts transitively on the set of nontrivial elements of $U$ and on the set of nontrivial irreducible characters of $U$. Fix a nontrivial element $g^{+}$ of $U$, which is then a transvection. Since $M^{+}$ is rational, ${}^{*}\!R^{G^{+}}_{M^{+}}(\chi^{+})$ is just the Harish-Chandra restriction of $\chi^{+}$ to $P$. The formula (4.9) now means that the fixed point subspace of $U$ in any module $V$ affording $\chi^{+}$ is zero. Hence the restriction of $V$ to $U$ affords all, and only, nontrivial irreducible characters of $U$, each with the same multiplicity
$e:=\chi^{+}(1)/(q^{N-1}-1).$
The sum of these characters is the regular character of $U$ minus $1_{U}$, hence takes value $-1$ at $g^{+}$. It follows that
(4.10) $\chi^{+}(g^{+})=-e=-\frac{\chi^{+}(1)}{q^{N-1}-1}.$
We will interpret this identity in terms of Green functions of $L^{+}=\mathrm{GL}_{2}(q^{p})$. According to [2, p. 191],
$R^{L^{+}}_{T^{+}_{1,1}}(1_{T^{+}_{1,1}})=1_{L^{+}}+\mathsf{St}_{L^{+}},~{}R^{L^{+}}_{T^{+}_{2}}(1_{T^{+}_{2}})=1_{L^{+}}-\mathsf{St}_{L^{+}}.$
Here, $T^{+}_{1,1}$ is a maximal torus in $L^{+}$ of order $(q^{p}-1)^{2}$ and $T^{+}_{2}$ is a maximal torus in $L^{+}$ of order $q^{2p}-1$. Viewed in $G^{+}$, they become maximal tori $T^{+}_{p,p}$ and $T^{+}_{2p}$, corresponding to permutations of type $(p,p)$ and $(2p)$ in the Weyl group $\mathsf{S}_{2p}$. Applying Lusztig induction to $G^{+}$, we get
(4.11) $R^{G^{+}}_{T^{+}_{p,p}}(1_{T^{+}_{p,p}})=R^{G^{+}}_{L^{+}}(1_{L^{+}})+R^{G^{+}}_{L^{+}}(\mathsf{St}_{L^{+}}),~{}R^{G^{+}}_{T^{+}_{2p}}(1_{T^{+}_{2p}})=R^{G^{+}}_{L^{+}}(1_{L^{+}})-R^{G^{+}}_{L^{+}}(\mathsf{St}_{L^{+}}).$
We now evaluate these formulae at the unipotent element $g^{+}\in G^{+}$, and use the fact that $R^{G^{+}}_{L^{\prime}}(\alpha)$ and $R^{G^{+}}_{L^{\prime}}(\theta\alpha)$ agree at the unipotent element $g^{+}$ for any linear character $\theta$ of a Levi subgroup $L^{\prime}$ of $G^{+}$, cf. [2, Proposition 10.1.2]. Also, each of $R^{G^{+}}_{L^{+}}(\theta\cdot 1_{L^{+}})$ and $R^{G^{+}}_{L^{+}}(\theta\cdot\mathsf{St}_{L^{+}})$ is, up to sign, an irreducible character $\chi^{+}$ corresponding to $L^{+}$, for a suitable choice of the linear character $\theta$. Hence (4.10) implies the following equalities for Green functions
$Q^{G^{+}}_{T^{+}_{p,p}}(g^{+})=-\frac{Q^{G^{+}}_{T^{+}_{p,p}}(1)}{q^{N-1}-1},~{}Q^{G^{+}}_{T^{+}_{2p}}(g^{+})=-\frac{Q^{G^{+}}_{T^{+}_{2p}}(1)}{q^{N-1}-1}.$
Since we have proved these equalities for all prime powers $q$, they also hold as identities for Green functions considered as polynomials in $q$. Applying the Ennola duality $q\mapsto-q$, cf. [21, §15.9], we now obtain
(4.12) $Q^{G}_{T_{p,p}}(g)=\frac{Q^{G}_{T_{p,p}}(1)}{q^{N-1}+1},~{}Q^{G}_{T_{2p}}(g)=\frac{Q^{G}_{T_{2p}}(1)}{q^{N-1}+1}$
for Green functions of $G=\mathrm{GU}_{2p}(q)$, where $T_{p,p}$ is a maximal torus of order $(q^{p}+1)^{2}$ and $T_{2p}$ is a maximal torus of order $q^{2p}-1$ of $G$.
We can take $T_{p,p}$ to be a maximal torus $T_{1,1}$ of $L=\mathrm{GU}_{2}(q^{p})$ and $T_{2p}=T$ to be a maximal torus $T_{2}$ of $L$. Inside $L$ we now have
$R^{L}_{T_{1,1}}(1_{T_{1,1}})=1_{L}-\mathsf{St}_{L},~{}R^{L}_{T_{2}}(1_{T_{2}})=1_{L}+\mathsf{St}_{L}.$
After Lusztig inducing to $G$, we obtain
$R^{G}_{T_{p,p}}(1_{T_{p,p}})=R^{G}_{L}(1_{L})-R^{G}_{L}(\mathsf{St}_{L}),~{}R^{G}_{T_{2p}}(1_{T_{2p}})=R^{G}_{L}(1_{L})+R^{G}_{L}(\mathsf{St}_{L}),$
and so
(4.13) $R^{G}_{L}(1_{L})=\bigl{(}R^{G}_{T_{p,p}}(1_{T_{p,p}})+R^{G}_{T_{2p}}(1_{T_{2p}})\bigr{)}/2,~{}R^{G}_{L}(\mathsf{St}_{L})=\bigl{(}R^{G}_{T_{p,p}}(1_{T_{p,p}})-R^{G}_{T_{2p}}(1_{T_{2p}})\bigr{)}/2.$
Note that each irreducible character $\chi$ of type (III) is, up to sign, $R^{G}_{L}(\theta\cdot 1_{L})$ or $R^{G}_{L}(\theta\cdot\mathsf{St}_{L})$ for a suitable linear character $\theta$ of $L$, and the values at $g$ of the latter characters are equal to the values at $g$ of $R^{G}_{L}(1_{L})$, respectively $R^{G}_{L}(\mathsf{St}_{L})$. Hence (4.12) and (4.13) show that (4.8) holds for every character of type (III). Finally, each irreducible character $\chi$ of of type (IV) is, up to sign, $R^{G}_{T_{2p}}(\theta)$ for a suitable linear character $\theta$ of $T_{2p}$. Hence, (4.12) shows that (4.8) holds for every character of type (IV), and the proof is complete. ∎
5. Thompson’s conjecture for $\mathrm{PSU}_{2p+1}(q)$
Let $p\geq 3$ be any prime, and let $q$ be any prime power (which may be coprime to $p$). In this section we show that Thompson’s conjecture holds for the simple group $\mathrm{PSU}_{2p+1}(q)$ as long as $p$ is sufficiently large.
We will work inside $G:=\mathrm{GU}_{2p+1}(q)$ and consider any semisimple element $x\in G$ which has both some fixed $\xi\in\mathbb{F}_{q^{p}}\smallsetminus\mathbb{F}_{q}$ and $1$ as eigenvalues. As in §4, this choice of $\xi$ implies that $x$ is regular semisimple, real, and of determinant $1$.
Theorem 4.
There is an explicit absolute constant $A>0$ such that the following statement holds when the prime $p$ is at least $A$. If $C$ denotes the conjugacy class of the image of the element $x$ described above in the simple group $S=\mathrm{PSU}_{2p+1}(q)$, then $S=C^{2}$. In fact, $(x^{\mathrm{SU}_{2p}(q)})^{2}$ contains all unipotent elements and all elements of support $\geq 2$.
The rest of the section is devoted to the proof of Theorem 4. By the main result of [3], it suffices to prove the result for $q\leq 7$. The chosen element $x$ has
$T:=\mathbf{C}_{G}(x)=T_{1}\times T_{2},\mbox{ where }T_{1}=\mathrm{GL}_{1}(q^{2p})\cong C_{q^{2p}-1}\mbox{ and }T_{2}=\mathrm{GU}_{1}(q)\cong C_{q+1}.$
In particular, $T\cap\mathrm{SU}_{2p+1}(q)$ has index $q+1$ in $T$, and hence the conjugacy class $x^{G}$ is a single $\mathrm{SU}_{2p+1}(q)$-conjugacy class. As $x$ is real, it suffices to show that, any non-central element $g\in\mathrm{SU}_{2p+1}(q)$ is a product of two $G$-conjugates of $x$, except possibly in the case $g=zh$ where $1\neq z\in\mathbf{Z}(\mathrm{SU}_{2p+1}(q))$ and $h$ is a transvection. By [9, Lemma 5.1], the statement holds in the case $g$ is semisimple. So we may assume that $g$ is not semisimple, and furthermore $g$ is a transvection if $\mathsf{supp}(g)=1$.
Note that the statement also follows in the case
(5.1) $g=\mathrm{diag}(g^{\prime},1)\in\mathrm{SU}_{2p}(q)\times\mathrm{SU}_{1}(q)\mbox{ with }g^{\prime}\in\mathrm{SU}_{2p}(q)\mbox{ but}$
$g^{\prime}\neq zh,\mbox{ with }1\neq z\in\mathbf{Z}(\mathrm{SU}_{2p}(q))\mbox{ and }h\mbox{ is a transvection}.$
Indeed, as $g$ is not semisimple, $g^{\prime}$ is not central in $\mathrm{SU}_{2p}(q)$. We may put $x$ in $\mathrm{SU}_{2p}(q)$, and then apply Theorem 3 to the element $g^{\prime}$. Therefore, in what follows we may assume that
$m=\mathsf{supp}(g)\geq 2.$
As in §4, we use the Frobenius formula in the form (4.1) and identify the dual group $G^{*}$ with $G$. Again, any non-central semisimple element $s\in G$ has the property that $\mathbf{C}_{G}(s)$ is a Levi subgroup $L$ of $G$. Fixing an embedding of $\overline{\mathbb{F}_{q}}^{\times}$ into $\mathbb{C}^{\times}$, we can identify $\mathbf{Z}(\mathbf{C}_{G}(s))$ with
$\mathrm{Hom}(\mathbf{C}_{G}(s)/[\mathbf{C}_{G}(s),\mathbf{C}_{G}(s)],\mathbb{C}^{\times}),$
and denote the linear character of $L=\mathbf{C}_{G}(s)$ corresponding to $s$ by $\hat{s}$. Then, any irreducible character $\chi$ in the Lusztig series labeled by $s$ can be written in the form (4.3), where $G$ and $L$ are understood to be defined as in this section. Assume now that $\chi(x)\neq 0$. As $x$ is regular semisimple, (4.4) implies that, up to conjugacy, $x\in L=\mathbf{C}_{G}(s)$. It follows that $s\in\mathbf{C}_{G}(x)=T$; in fact, $s\in T\smallsetminus\mathbf{Z}(G)$. Now using the primality of $p$, we see that the characters $\chi\in\mathrm{Irr}(G)$ that are non-vanishing at $x$ are divided into the following types:
  1. (I)
    Characters $\chi$ in (4.3) with $s\in T_{11}\times T_{2}<T$, where $T_{11}$ is the unique subgroup $C_{q+1}$ of $T_{1}$.
  2. (II)
    Characters $\chi$ in (4.3) with $s\in T_{12}\times T_{2}\smallsetminus(T_{11}\times T_{2})$, where $T_{12}$ is the unique subgroup $C_{q^{2}-1}$ of $T_{1}$ (in particular, $q>2$). For any such $s$, $\mathbf{C}_{G}(s)\cong\mathrm{GL}_{p}(q^{2})\times\mathrm{GU}_{1}(q)$.
  3. (III)
    Characters $\chi$ in (4.3) with $s\in T_{13}\times T_{2}\smallsetminus(T_{11}\times T_{2})$, where $T_{13}$ is the unique subgroup $C_{q^{p}+1}$ of $T_{1}$. For any such $s$, $\mathbf{C}_{G}(s)\cong\mathrm{GU}_{2}(q^{p})\times\mathrm{GU}_{1}(q)$.
  4. (IV)
    Characters $\chi$ in (4.3) with $s\in T$ not belonging to the union of $T_{12}\times T_{2}$ and $T_{13}\times T_{2}$. Any such $s$ is regular semisimple, with $\mathbf{C}_{G}(s)=T$.
The Frobenius formula includes $q+1$ linear characters which all take value $1$ at any element in $\mathrm{SU}_{2p+1}(q)$. Hence it suffices to show
(5.2) |*$\Bigm{|}\sum\nolimits^{*}\frac{|\chi(x)|^{2}\chi(g)}{\chi(1)}\Bigm{|}<q+1,$
where, again, we sum only over $\chi$ with $\chi(1)>1$, $\chi(x)\neq 0$, and $\chi(g)\not\in[0,\infty)$. We again refer to the sum (5.2) as $\Sigma^{*}$. We will frequently use the character bound (4.5), with
$N:=2p+1.$
Proposition 5.1.
There is an absolute constant $N_{1}>0$ such that when $N\geq N_{1}$, the total contribution of characters of type (I) to $\Sigma^{*}$ has absolute value $<2(q+1)/3$.
Proof.
(a) The elements $s$ that label the characters $\chi$ of type (I) are
$s=\mathrm{diag}(aI_{N-1},b)$
where $a,b\in\mu_{q+1}\subseteq\mathbb{F}_{q^{2}}^{\times}$.
If $a=b$, then $\chi$ is a unipotent character $\chi^{\alpha}$ times a linear character. Since $x$ is regular semisimple with centralizer $T\cong C_{q^{N-1}-1}\times C_{q+1}$, $\chi(x)\neq 0$ only when the unipotent character $\chi^{\alpha}$ is labeled by one of the $N-1$ partitions $\alpha=(N)$, $(1^{N})$, or $(N-k,2,1^{N-2-k})$ with $22\leq k\leq N-$ 2, in which case
(5.3) $|\chi(x)|=1,$
cf. [13, Corollary 3.1.2].
Next we show that (5.3) also holds for the $\chi$ with $a\neq b$ and $\chi(x)\neq 0$. Indeed, we will use (4.4) and look at all $y\in G$ where $yxy^{-1}\in L=\mathbf{C}_{G}(s)=\mathrm{GU}_{N-1}(q)\times\mathrm{GU}_{1}(q)\ni x$. This implies that $y$ fixes the unique $\mathbb{F}_{q^{2}}$-line fixed by $x$ and hence $y\in L$. For such a $y$, the $\mathrm{GU}_{N-1}(q)$-component of $yxy^{-1}$ is a regular semisimple element of $\mathrm{GU}_{N-1}(q)$ with centralizer $C_{q^{N-1}-1}$, and by [13, Corollary 3.1.2], $\chi(x)\neq 0$ implies that the $\mathrm{GU}_{N-1}$-component of $\psi$ is one of the $N-1$ hook unipotent characters, each taking value ±$\pm 1$ at $yxy^{-1}$. As $\mathsf{St}_{G}(x)=\pm 1$, (4.4) implies (5.3).
(b) First we consider the $\chi$ with
$j:=\mathfrak{l}(\chi)\geq\sqrt{N/13}.$
Choosing $N_{1}$ sufficiently large, when $N\geq N_{1}$, by [7, Theorem 1.2] we have $\chi(1)\geq q^{c_{1}N^{3/2}}$ for some absolute constant $c_{1}>0$. So, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$2N(q+1)^{2}q^{-c_{1}\sigma\sqrt{N}}<\frac{q+1}{12}$
when $N_{1}$ is large enough.
Now we look at the $\chi$ with $1\leq j\leq\sqrt{N/13}$. Suppose that
$m=\mathsf{supp}(g)\geq\sqrt{N/13}.$
Since $\chi(1)\geq q^{N/2}$, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$2\sqrt{N/13}(q+1)^{2}q^{-\sigma\sqrt{N/52}}<\frac{q+1}{3}$
when $N_{1}$ is large enough, yielding the claim in this case.
(c) In the remaining case, we have $2\leq m\leq\sqrt{N/13}$ and $j\leq\sqrt{N/13}$. Still, for each $j$ we have at most $(q+1)^{2}$ characters $\chi$ to consider, one for each pair $(a,b)$. Let $\delta_{a}$ denote the linear character $\hat{t}$ for $t=aI_{N}$. Then we can write $\chi=\delta_{a}\chi^{\prime}$, where $\chi^{\prime}$ now has $\mathfrak{l}^{*}(\chi^{\prime})=j$ by [7, Theorem 3.9], and $\delta_{a}(g)=1$ as $g\in\mathrm{SU}_{N}(q)$. By Theorem 2,
$\frac{\chi(g)}{\chi(1)}=\frac{\chi^{\prime}(g)}{\chi^{\prime}(1)}=\epsilon_{\chi}+(-q)^{-mj}\gamma_{\lambda},$
where $|\epsilon_{\chi}|<q^{-N/2}$ and $\gamma_{\lambda}=\chi^{\prime}(\lambda I_{A})/\chi^{\prime}(1)$ for the primary eigenvalue $\lambda$ of $g$.
If $m\geq 3$, then
$(q+1)^{2}\Bigm{|}\sum_{j\geq 1}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq(q+1)^{2}\sum^{\infty}_{j\geq 1}q^{-3j}=(q+1)^{2}/(q^{3}-1)\leq\frac{3(q+1)}{7}.$
If $m=2$ but $q\geq 3$, then
$(q+1)^{2}\Bigm{|}\sum_{j\geq 1}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq(q+1)^{2}\sum^{\infty}_{j\geq 1}q^{-2j}=(q+1)^{2}/(q^{2}-1)\leq\frac{q+1}{2}.$
Suppose $m=2=q$. Recall that $g$ is a non-semisimple element of support $2$ by our convention. If $\lambda=1$, then $g$ has the shape (5.1), and so (4.1) is already established. So assume $\lambda\neq 1$. Since $q=2$, we have $\lambda=\omega$, a fixed cubic root of unity in $\mathbb{F}_{4}$. For the three choices $(a,b)=(1,1)$, $(\omega,\omega)$, $(\omega^{2},\omega^{2})$, $\chi^{\prime}$ is the same unipotent character, for which $\gamma_{\lambda}=1$. For the three choices $(a,b)=(1,\omega)$, $(\omega,\omega^{2})$, $(\omega^{2},1)$, $\chi^{\prime}$ is the same character corresponding to $s=\mathrm{diag}(I_{N-1},\omega)\notin\mathrm{SU}_{N}(q)$, for which $\gamma_{\lambda}=\zeta$, a fixed cubic root of unity in $\mathbb{C}$, see [20, Proposition 4.5]. For the three choices $(a,b)=(\omega,1)$, $(\omega^{2},\omega)$, $(1,\omega^{2})$, $\chi^{\prime}$ is the complex conjugate of the preceding character (and corresponding to $s=\mathrm{diag}(I_{N-1},\omega^{2})$), for which $\gamma_{\lambda}=\bar{\zeta}$. Hence the sum of the terms $(-q)^{mj}\gamma_{\lambda}$ over these nine characters for any fixed $j$ is 00.
Hence, choosing $N_{1}$ large enough, for all possible $(m,q)$, we have
$(q+1)^{2}\Bigm{|}\sum_{1\leq j\leq\sqrt{N/13}}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq(q+1)\Bigl{(}\frac{1}{2}+\frac{1}{24}\Bigr{)}$
and
$(q+1)^{2}\Bigm{|}\sum_{1\leq j\leq\sqrt{N/13}}\epsilon_{j}\Bigm{|}\leq(q+1)^{2}q^{-N/2}\sqrt{N/13}\leq\frac{q+1}{24}.$
It follows that
$\Bigm{|}\sum_{1\leq j=\mathfrak{l}(\chi)\leq\sqrt{N/13}}\frac{|\chi(x)|^{2}\chi_{j}(g)}{\chi_{j}(1)}\Bigm{|}\leq(q+1)\Bigl{(}\frac{1}{2}+\frac{1}{12}\Bigr{)}=\frac{7(q+1)}{12},$
and so we are done in this case as well. ∎
Lemma 5.2.
There is an absolute constant $N_{2}>0$ such that when $N\geq N_{2}$, the total contribution of characters of type (II) to $\Sigma^{*}$ has absolute value $<1/6$.
Proof.
Each of these characters can be written in the form (4.3), where $L=\mathrm{GL}_{p}(q^{2})\times\mathrm{GU}_{1}(q)$ is unique up to conjugacy in $G$. For the element $x$, we can again use the general bound (4.6). There are fewer than $(q+1)q^{2}<400$ choices for the element $s$ (up to conjugacy). Furthermore, by (4.4), the condition $\chi(x)\neq 0$ implies that (after conjugating $x$ suitably) $x\in L$ and $\psi(x)\neq 0$. Viewed as an element in $L$, the $\mathrm{GL}_{p}(q^{2})$-component of $x$ is again regular semisimple with centralizer $C_{q^{2p}-1}$. Hence the $\mathrm{GL}_{p}(q^{2})$-component of the unipotent character $\psi$ is labeled by one of the $p$ hook partitions of $p=(N-1)/2$. Thus there are fewer than $200N$ characters of type (II) with $\chi(x)\neq 0$. Each of them has degree
$\chi(1)\geq[G:L]_{q^{\prime}}=(q^{3}+1)(q^{5}+1)\ldots(q^{N-2}+1)(q^{N}+1)>q^{(N+1)^{2}/4-1}.$
Hence, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$200N^{3}q^{-\sigma(N/4+1/2-3/4N)}<\frac{1}{6}$
when $N\geq N_{2}$ and $N_{2}$ is large enough. ∎
To complete the proof of (5.2) and of Theorem 4, it remains to prove:
Lemma 5.3.
There is an absolute constant $N_{3}>0$ such that when $N\geq N_{3}$, the total contribution of characters of types (III) and (IV) to $\Sigma^{*}$ has absolute value $<1/6$.
Proof.
Since $\mathsf{supp}(g)\geq 2$, by Lemma 3.1 we have
$|\chi(g)|\leq\sqrt{|\mathbf{C}_{G}(g)|}\leq q^{N^{2}/2-1.7N+2.5}$
for any $\chi$; and [14, Corollary 7.6] still gives the bound (4.6) for $|\chi(x)|$.
Each $\chi$ of type (III) can be written in the form (4.3), where $L=\mathrm{GU}_{2}(q^{p})\times\mathrm{GU}_{1}(q)$ is unique up to conjugacy in $G$, and
$\chi(1)$ $\geq[G:L]_{q^{\prime}}$
$=(q^{2}-1)(q^{3}+1)\ldots(q^{p-1}-1)\cdot(q^{p+1}-1)(q^{p+2}+1)\ldots(q^{2p-1}+1)(q^{2p+1}+1)$
$>q^{N^{2}/2-N-1/2}$
by [11, Lemma 4.1(iii)]. There are fewer than $(q+1)q^{(N-1)/2}$ choices for $s$, and two choices for the unipotent character $\psi$. Thus there are fewer than $2(q+1)q^{(N-1)/2}<q^{(N+5)/2}$ choices for $\chi$ of type (III). Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$N^{2}q^{(N+5)/2}q^{-0.7N+3}\leq N^{2}2^{5.5-0.2N}<\frac{1}{12}$
when $N\geq N_{3}$ and $N_{3}$ is large enough.
Each $\chi$ of type (IV) can be written in the form (4.3), where $L=\mathrm{GL}_{1}(q^{2p})\times\mathrm{GU}_{1}(q)$ is unique up to conjugacy in $G$, and
$\chi(1)\geq[G:L]_{q^{\prime}}=(q^{2}-1)(q^{3}+1)\ldots(q^{2p-1}+1)(q^{2p+1}+1)>q^{(N^{2}-N)/2-1}.$
There are fewer than $(q+1)q^{N-1}<q^{N+1}$ choices for $s$, and one choice for the unipotent character $\psi$. Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$N^{2}q^{N+1}q^{-1.2N+3.5}\leq N^{2}2^{4.5-0.2N}<\frac{1}{12}$
when $N\geq N_{3}$ and $N_{3}$ is large enough. ∎
6. Thompson’s conjecture for $\mathrm{PSU}_{2p+2r+\kappa}(q)$
Let $p,r\geq 3$ be primes, $q$ be any prime power, and $\kappa\in\{0,1\}$. In this section we show that Thompson’s conjecture holds for the simple group $\mathrm{PSU}_{2p+2r+\kappa}(q)$ as long as $p$ is sufficiently large, and $r$ is suitably chosen.
We work inside the general unitary group $G=\mathrm{GU}_{N}(q)$ where
$N:=2p+2r+\kappa,$
and consider any semisimple element
$x=\mathrm{diag}(x_{2p},x_{2r},x_{\kappa})\in G,$
such that $x_{2p}\in\mathrm{GU}_{2p}(q)$ has some fixed $\xi\in\mathbb{F}_{q^{p}}\smallsetminus\mathbb{F}_{q}$ as an eigenvalue, $x_{2r}\in\mathrm{GU}_{2r}(q)$ has some fixed $\xi^{\prime}\in\mathbb{F}_{q^{r}}\smallsetminus\mathbb{F}_{q}$ as an eigenvalue. Furthermore, $x_{\kappa}$ occurs only when $\kappa=1$, in which case $x_{\kappa}=I_{1}$. This ensures that $x$ is regular semisimple, real, and of determinant $1$.
Theorem 5.
There exist explicit absolute constants $A_{3}>A_{2}>A_{1}>0$ such that the following statements hold for any prime $p\geq A_{3}$, any prime $r$ in the interval $[A_{1},A_{2}]$, any prime power $q\leq 7$, and any $\kappa\in\{0,1\}$.
  1. (i)
    The interval $[A_{1},A_{2}]$ contains prime integers that are congruent to any fixed congruence class $\ell$ modulo $120$ with $\gcd(\ell,120)=1$.
  2. (ii)
    If $C$ denotes the conjugacy class of the image of the element $x$ described above in the simple group $S:=\mathrm{PSU}_{2p+2r+\kappa}(q)$, then $S=C^{2}$. In fact, $(x^{\mathrm{SU}_{2p+2r+\kappa}(q)})^{2}$ contains all unipotent elements and all elements of support $\geq 2$.
Certainly, condition 5(i) holds when both $A_{1}$ and $A_{2}-A_{1}$ are large enough. The rest of the section is devoted to the proof of Theorem 5(ii).
The chosen element $x$ has
$T:=\mathbf{C}_{G}(x)=T_{1}\times T_{2}\times T_{3},\mbox{ where }T_{1}=\mathrm{GL}_{1}(q^{2p})\cong C_{q^{2p}-1}\mbox{ and }T_{2}=\mathrm{GL}_{1}(q^{2r})\cong C_{q^{2r}-1};$
furthermore, $T_{3}$ occurs only when $\kappa=1$, in which case $T_{3}=\mathrm{GU}_{1}(q)$. In particular, $T\cap\mathrm{SU}_{N}(q)$ has index $q+1$ in $T$, and hence the conjugacy class $x^{G}$ is a single $\mathrm{SU}_{N}(q)$-conjugacy class. As $x$ is real, it suffices to show that, any non-central element $g\in\mathrm{SU}_{N}(q)$ is a product of two $G$-conjugates of $x$, except possibly in the case $g=zh$ where $1\neq z\in\mathbf{Z}(\mathrm{SU}_{N}(q))$ and $h$ is a transvection. Using [9, Lemma 5.1] as before, we may assume that $g$ is not semisimple, and furthermore $g$ is a transvection if $\mathsf{supp}(g)=1$.
Note that the statement also follows in the case
(6.1) $g=\mathrm{diag}(g^{\prime},I_{2r+\kappa})\in\mathrm{SU}_{2p}(q)\times\mathrm{SU}_{2r+\kappa}(q)\mbox{ with }g^{\prime}\in\mathrm{SU}_{2p}(q)$
$\mbox{but }g^{\prime}\neq zh,\mbox{ with }1\neq z\in\mathbf{Z}(\mathrm{SU}_{2p}(q))\mbox{ and }h\mbox{ is a transvection}.$
Indeed, as $g$ is not semisimple, $g^{\prime}$ is not central in $\mathrm{SU}_{2p}(q)$. We may put $x$ in $\mathrm{SU}_{2p}(q)$, and then apply Theorem 3 to the element $g^{\prime}$. Therefore, in what follows we may assume that
$m=\mathsf{supp}(g)\geq 2.$
As before, we need to show (4.1). We identify the dual group $G^{*}$ with $G$ and express any irreducible character $\chi$ in the Lusztig series labeled by a semisimple $s\in G$ such that $\mathbf{C}_{G}(s)=L$ in the form (4.3). Assume now that $\chi(x)\neq 0$. As $x$ is regular semisimple, (4.4) implies that, up to conjugacy, $x\in L=\mathbf{C}_{G}(s)$. It follows that $s\in\mathbf{C}_{G}(x)=T$; write
$s=\mathrm{diag}(s_{2p},s_{2r},s_{\kappa}),$
where $s_{\kappa}$ occurs only when $\kappa=1$. Now using the primality of $p$ and of $r$, we see that the characters $\chi\in\mathrm{Irr}(G)$ that are non-vanishing at $x$ are divided into the following types:
  1. (I)
    Characters $\chi$ in (4.3) with $s_{2p}\in T_{11}$, where $T_{11}$ is the unique subgroup $C_{q+1}$ of $T_{1}$.
  2. (II)
    Characters $\chi$ in (4.3) with $s_{2p}\in T_{12}\smallsetminus T_{11}$, where $T_{12}$ is the unique subgroup $C_{q^{2}-1}$ of $T_{1}$ (in particular, $q>2$). Either
    (6.2) $\mathbf{C}_{G}(s)=\mathrm{GL}_{p}(q^{2})\times\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(s)$
    or
    (6.3) $\mathbf{C}_{G}(s)=\mathrm{GL}_{p+r}(q^{2})\times T_{3},$
    depending on whether the set of eigenvalues of $s_{2r}$ is different from or the same as the set of eigenvalues of $s_{2p}$.
  3. (III)
    Characters $\chi$ in (4.3) with $s_{2p}\in T_{13}\smallsetminus T_{11}$, where $T_{13}$ is the unique subgroup $C_{q^{p}+1}$ of $T_{1}$. For any such $s$,
    (6.4) $\mathbf{C}_{G}(s)=\mathrm{GU}_{2}(q^{p})\times\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(s).$
  4. (IV)
    Characters $\chi$ in (4.3) with $s_{2p}\in T_{1}\smallsetminus(T_{12}\cup T_{13})$. For any such $s$,
    (6.5) $\mathbf{C}_{G}(s)=T_{1}\times\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(s).$
Here $\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(s)$ is short for $\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(\mathrm{diag}(s_{2r},s_{\kappa}))$.
We note that there is an explicit function $f(A_{2})$ (which can be taken to be the class number of $\mathsf{S}_{2A_{2}+1}$) such that the Levi subgroup $\mathbf{C}_{\mathrm{GU}_{2r+\kappa}(q)}(s)$ of $\mathrm{GU}_{2r+\kappa}(q)$ has at most $f(A_{2})$ unipotent characters. We denote by $T_{21}$ the unique subgroup $C_{q+1}$ of $T_{2}$.
The Frobenius sum includes $q+1$ linear characters which all take value $1$ at any element in $\mathrm{SU}_{N}(q)$. Hence it suffices to show
(6.6) |*$\Bigm{|}\sum\nolimits^{*}\frac{|\chi(x)|^{2}\chi(g)}{\chi(1)}\Bigm{|}<q+1,$
where we again refer to the summation in (6.6), which is restricted to $\chi\in\mathrm{Irr}(G)$ such that $\chi(1)>1$, $\chi(x)\neq 0$, and $\chi(g)\not\in[0,\infty)$, as $\Sigma^{*}$ and frequently use the character bound (4.5). For the element $x$, we have the bound
(6.7) $|\chi(x)|\leq 4pr$
which follows from [14, Corollary 7.6].
Proposition 6.1.
The following statement holds when $A_{1}$ is large enough. For each $A_{2}>A_{1}$, there is an absolute constant $N_{1}>0$ (depending on $A_{2}$) such that when $A_{1}\leq r\leq A_{2}$ and $N\geq N_{1}$, the total contribution of characters of type (I) to $\Sigma^{*}$ has absolute value $<7(q+1)/8$.
Proof.
(a) We will choose $N\geq N_{1}>8A_{2}+1$ so that $p>3r$. The elements $s$ that label the characters $\chi$ of type (I) are
$s=\mathrm{diag}(aI_{2p},s_{2r},cI_{\kappa})$
where $a$ and, if $\kappa=1$, also $c$, belong to $\mu_{q+1}\subseteq\mathbb{F}_{q^{2}}^{\times}$, and $s_{2r}\in\mathrm{GL}_{1}(q^{2r})$.
First we show that
(6.8) $|\chi(x)|=1$
if $\chi(x)\neq 0$ and $s_{2r}\in T_{21}$, that is, $s_{2r}=bI_{2r}$ with $b\in\mu_{q+1}$.
Suppose that $a=b$ and furthermore $a=c$ if $\kappa=1$. Then $\chi$ is a unipotent character $\chi^{\alpha}$ of $G$ times a linear character. Recall that $x$ is regular semisimple with centralizer $T\cong C_{q^{2p}-1}\times C_{q^{2r}-1}\times C_{q+1}$ if $\kappa=1$ and $T\cong C_{q^{2p}-1}\times C_{q^{2r}-1}$ if $\kappa=0$. Hence $\chi(x)\neq 0$ only when the unipotent character $\chi^{\alpha}$ is labeled by one of the partitions prescribed by Proposition 2.2 or Proposition 2.3 for $\kappa=0$ or $\kappa=1$ respectively, and in either case, (6.8) holds.
Next suppose that $\kappa=1$ and $a=b\neq c$. We now use (4.4) and look at all $y\in G$ where $yxy^{-1}\in L=\mathbf{C}_{G}(s)=\mathrm{GU}_{N-1}(q)\times\mathrm{GU}_{1}(q)\ni x$. This implies that $y$ fixes the unique $\mathbb{F}_{q^{2}}$-line fixed by $x$ in $\mathbb{F}_{q^{2}}^{N}$, and hence $y\in L$. For such an $y$, the $\mathrm{GU}_{N-1}(q)$-component of $yxy^{-1}$ is a regular semisimple element of $L$ with centralizer $C_{q^{2p}-1}\times C_{q^{2r}-1}$. Now $\chi(x)\neq 0$ implies that the $\mathrm{GU}_{N-1}$-component of $\psi$ is labeled by one of the partitions prescribed by Proposition 2.2 and takes value ±$\pm 1$ at $yxy^{-1}$. As $\mathsf{St}_{G}(x)=\pm 1$, (4.4) implies (6.8).
Next suppose that $\kappa=0$ and $a\neq b$. We again use (4.4) and look at all $y\in G$ where $yxy^{-1}\in L=\mathbf{C}_{G}(s)=\mathrm{GU}_{2p}(q)\times\mathrm{GU}_{2r}(q)\ni x$. Since $x$ fixes a unique non-degenerate subspace $M$ of dimension $2r$ in $\mathbb{F}_{q^{2}}^{N}$, this implies that $y$ fixes $M$ and hence $y\in L$. For such a $y$, the $\mathrm{GU}_{2p}(q)$-component of $yxy^{-1}$ is a regular semisimple element of $\mathrm{GU}_{2p}(q)$ with centralizer $C_{q^{2p}-1}$, and the $\mathrm{GU}_{2r}(q)$-component of $yxy^{-1}$ is a regular semisimple element of $\mathrm{GU}_{2r}(q)$ with centralizer $C_{q^{2r}-1}$. Now $\chi(x)\neq 0$ implies that the $\mathrm{GU}_{2p}(q)$-component of $\psi$ is labeled by one of the $2p$ hook partitions of $2p$, and the $\mathrm{GU}_{2r}(q)$-component of $\psi$ is labeled by one of the $2r$ hook partitions of $2r$, each taking value ±$\pm 1$ at the corresponding component of $yxy^{-1}$. As $\mathsf{St}_{G}(x)=\pm 1$, (4.4) implies (6.8).
The arguments in the preceding paragraph also apply to each of the remaining possibilities when $a\neq b$ and $\kappa=1$:
$\bullet$ $c=a\neq b$;
$\bullet$ $c=b\neq a$; and
$\bullet$ $c\notin\{a,b\}$.
In each of the above possibilities with $s_{2r}\in T_{21}$, our analysis also shows that the number of $\chi$ with $\chi(x)\neq 0$ is at most $(q+1)^{3}(2p)(2r)\leq 2^{10}NA_{2}$.
If $s_{2r}\notin T_{21}$, then $|\chi(x)|\leq 4pr<2NA_{2}$ by (6.8), and the number of $\chi$ with $\chi(x)\neq 0$ is at most $(q+1)^{2}q^{2r}(2p+\kappa)f(A_{2})<2^{6}Nq^{2A_{2}}f(A_{2})$ (here, $2p+\kappa$ accounts for the hook unipotent characters of $\mathrm{GU}_{2p}(q)$ or $\mathrm{GU}_{2p+\kappa}(q)$).
Thus the total number of $\chi$ of type (I) with $\chi(x)\neq 0$ is at most $65Nq^{2A_{2}}f(A_{2})$ when $A_{2}$ is large enough.
(b) First we consider the $\chi$ with
$j:=\mathfrak{l}(\chi)\geq\sqrt{N/13}.$
Choosing $N_{1}$ sufficiently large, when $N\geq N_{1}$, by [7, Theorem 1.2] we have $\chi(1)\geq q^{c_{1}N^{3/2}}$ for some absolute constant $c_{1}>0$. So, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$260N^{3}A_{2}^{2}f(A_{2})q^{2A_{2}}q^{-c_{1}\sigma\sqrt{N}}<\frac{q+1}{12}$
when $N_{1}$ is large enough compared to $A_{2}$.
Now we look at the $\chi$ with $1\leq j\leq\sqrt{N/13}$. Suppose that
$m=\mathsf{supp}(g)\geq\sqrt{N/13}.$
Since $\chi(1)\geq q^{N/2}$, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$260N^{3}A_{2}^{2}f(A_{2})q^{2A_{2}}q^{-\sigma\sqrt{N/52}}<\frac{q+1}{2}$
when $N_{1}$ is large enough compared to $A_{2}$, yielding the claim in this case.
(c) In the remaining case, we have $2\leq m\leq\sqrt{N/13}$ and $j\leq\sqrt{N/13}$. Since $p\geq r+1$ and $s_{2p}=aI_{2p}$, $a$ is the primary eigenvalue of $s$. Let $\delta_{a}$ denote the linear character $\hat{t}$ of $G$ for $t=aI_{N}$. Then we can write
$\chi=\delta_{a}\chi^{\prime},$
where $\chi^{\prime}$ now has $\mathfrak{l}^{*}(\chi^{\prime})=j$ by [7, Theorem 3.9], and $\delta_{a}(g)=1$ as $g\in\mathrm{SU}_{N}(q)$. By Theorem 2,
(6.9) $\frac{\chi(g)}{\chi(1)}=\frac{\chi^{\prime}(g)}{\chi^{\prime}(1)}=\epsilon_{\chi}+(-q)^{-mj}\gamma_{\lambda},$
where $|\epsilon_{\chi}|<q^{-N/2}$ and $\gamma_{\lambda}=\chi^{\prime}(\lambda I_{A})/\chi^{\prime}(1)$ for the primary eigenvalue $\lambda$ of $g$.
Note that the total contribution of the error terms $\epsilon_{\chi}$ in (6.9) is at most
(6.10) $260N^{3}A_{2}^{2}f(A_{2})q^{2A_{2}}q^{-N/2}<\frac{q+1}{12}$
when $N_{1}$ is large enough compared to $A_{2}$. So we need to bound the total contribution of the main term $(-q)^{-mj}\gamma_{\lambda}$ in (6.9); call it the total major contribution.
First we bound the major contribution coming from those $\chi$ where either $s_{2r}\notin T_{21}$ or $s_{2r}=bI_{2r}$ with $a\neq b\in\mu_{q+1}$. Then $a$ is not an eigenvalue of $s_{2r}$, and this already implies that
$j\geq 2r.$
By [7, Theorem 3.9], $j$ is determined by the partition $\alpha$, of $2p$ if $\kappa=0$ or if $\kappa=1$ but $c\neq a$, and of $12p+$ 1 if $\kappa=1$ and $a=c$, where $\alpha$ labels the corresponding component $\psi^{\alpha}$ of $\psi$ in the subgroup $\mathrm{GU}_{2p}(q)$, respectively $\mathrm{GU}_{2p+1}(q)$, of $L=\mathbf{C}_{G}(s)$, cf. the expression (4.3) for $\chi$. In the former case, the unipotent character $\psi^{\alpha}$ is nonzero at the regular semisimple $x_{2p}$, whence $\alpha$ is a hook partition of $2p$ and uniquely determined by $j$. In the latter case, $\psi^{\alpha}$ is nonzero at $\mathrm{diag}(x_{2p},x_{\kappa})$ of type $(2p,1)$. In this case $\alpha$ is one of $2p$ partitions of $12p+$ 1, see [13, Corollary 3.1.2], and is again uniquely determined by $j$. We have therefore shown that, for each choice of $s$ with $s_{2r}\notin T_{21}$, there is a unique $\chi$ of level $j$. As in (a), $\psi^{\alpha}$ takes value ±$\pm 1$ at the corresponding component of $x$. Inducing the $\mathbf{C}_{\mathrm{GU}_{2r}(q)}(s_{2r})$-component of $\psi$ to $\mathrm{GU}_{2r}(q)$ and using the fact that $x_{2r}$ is regular semisimple with centralizer $C_{q^{2r}-1}$ in $\mathrm{GU}_{2r}(q)$, by [21, Corollary 7.6] we have $|\chi(x)|\leq 2r$ using (4.4). Also, there are at most $(q+1)^{2}q^{2r}$ choices for $s$. Hence, the total major contribution coming from these $\chi$ is at most
$4r^{2}(q+1)^{2}q^{2r}\Bigm{|}\sum_{j\geq 2r}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq 4r^{2}(q+1)^{2}q^{2r}\sum^{\infty}_{j\geq 2r}q^{-2j}=\frac{4r^{2}(q+1)^{2}}{q^{2r-2}(q^{2}-1)}\leq(q+1)/12$
when $r\geq A_{1}$ and $A_{1}$ is large enough.
(d) It remains to bound the major contribution coming from those $\chi$ with
(6.11) $s=\mathrm{diag}(aI_{2p+2r},cI_{\kappa})$
where $a$ and, if $\kappa=1$, also $c$, belong to $\mu_{q+1}$. Here, $\mathbf{C}_{G}(s)=\mathrm{GU}_{2p+2r}(q)\times T_{3}$ if $\kappa=0$ or if $\kappa=1$ but $a\neq c$, and $\mathbf{C}_{G}(s)=\mathrm{GU}_{2p+2r+1}(q)$ if $\kappa=1$ and $a=c$. By (6.8) we have $|\chi(x)|=1$. By [7, Theorem 3.9], $j$ is determined by the partition $\alpha$ of $2p+2r$, respectively $12p+2r+$ 1, where $\alpha$ labels the corresponding component $\psi^{\alpha}$ of $\psi$ in the subgroup $\mathrm{GU}_{2p+2r}(q)$, respectively $\mathrm{GU}_{2p+2r+1}(q)$, of $L=\mathbf{C}_{G}(s)$. In the former case, the unipotent character $\psi^{\alpha}$ is nonzero at the regular semisimple element $\mathrm{diag}(x_{2p},x_{2r})$ of type $(2p,2r)$. By Proposition 2.2, for a given $j$ the number of $\alpha$ is $\leq 1$ if $11\leq j\leq 2r-$ 1 and at most $12r^{2}+r+$ 1 in general. In the latter case, $\psi^{\alpha}$ is nonzero at the regular semisimple $\mathrm{diag}(x_{2p},x_{2r},x_{\kappa})$ of type $(2p,2r,1)$. By Proposition 2.3, for a given $j$ the number of $\alpha$ is $\leq 1$ if $11\leq j\leq 2r-$ 1 and at most $(2r+1)^{2}$ in general.
Now, the total major contribution coming from those $\chi$ subject to (6.11) and of level $2r\leq j\leq\sqrt{N/13}$ is at most
$(2r+1)^{2}(q+1)^{2}\Bigm{|}\sum_{j\geq 2r}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}$ $\leq(2r+1)^{2}(q+1)^{2}\sum^{\infty}_{j\geq 2r}q^{-2j}$
$=\frac{(2r+1)^{2}(q+1)^{2}}{q^{2r-2}(q^{2}-1)}$
$\leq\frac{q+1}{8}$
when $r\geq A_{1}$ and $A_{1}$ is large enough.
The total major contribution coming from those $\chi$ subject to (6.11) and $11\leq j\leq 2r-$ 1 is at most
$(q+1)^{1+\kappa}\Bigm{|}\sum_{1\leq j<2r}(-q)^{-mj}\gamma_{\lambda}\Bigm{|}\leq(q+1)^{1+\kappa}\sum^{\infty}_{j\geq 1}q^{-mj}\leq\frac{(q+1)^{1+\kappa}}{q^{m-2}(q^{2}-1)}\leq\frac{q+1}{2}$
if $\kappa=0$, or if $\kappa=1$ but $m\geq 3$, or if $(\kappa,m)=(1,2)$ but $q\geq 3$.
Suppose $\kappa=1$ and $m=2=q$. Recall that $g$ is a non-semisimple element of support $2$ with primary eigenvalue $\lambda$ by our convention. If $\lambda=1$, then $g$ has the shape (6.1) and so (4.1) is already established. So assume that $\lambda\neq 1$. Since $q=2$, we have $\lambda=\omega$, a fixed cubic root of unity in $\mathbb{F}_{4}$. For the three choices $(a,c)=(1,1)$, $(\omega,\omega)$, $(\omega^{2},\omega^{2})$, $\chi^{\prime}$ is the same unipotent character, for which $\gamma_{\lambda}=1$. For the three choices $(a,c)=(1,\omega)$, $(\omega,\omega^{2})$, $(\omega^{2},1)$, $\chi^{\prime}$ is the same character corresponding to $s=\mathrm{diag}(I_{N-1},\omega)\notin\mathrm{SU}_{N}(q)$, for which $\gamma_{\lambda}=\zeta$, a fixed cubic root of unity in $\mathbb{C}$, see [20, Proposition 4.5]. For the three choices $(a,b)=(\omega,1)$, $(\omega^{2},\omega)$, $(1,\omega^{2})$, $\chi^{\prime}$ is the complex conjugate of the preceding character (and corresponding to $s=\mathrm{diag}(I_{N-1},\omega^{2})$), for which $\gamma_{\lambda}=\bar{\zeta}$. Hence the sum of the terms $(-q)^{mj}\gamma_{\lambda}$ over these nine characters for any fixed $j$ is 00.
Hence, in all cases, choosing $A_{1}$ large enough, and $N_{1}$ large enough compared to $A_{2}$, from the characters $\chi$ subject to (6.11) of level $\leq\sqrt{N/13}$ the major terms contribute at most $85(q+1)/$ 8. It follows that
$\Bigm{|}\sum_{1\leq j=\mathfrak{l}(\chi)}\frac{|\chi(x)|^{2}\chi_{j}(g)}{\chi_{j}(1)}\Bigm{|}\leq(q+1)\Bigl{(}\frac{5}{8}+\frac{1}{12}+\frac{1}{12}+\frac{1}{12}\Bigr{)}=\frac{7(q+1)}{8},$
and so we are done. ∎
Lemma 6.2.
For any constant $A_{2}\geq 3$, there is an absolute constant $N_{2}>0$ (depending on $A_{2}$) such that when $r\leq A_{2}$ and $N\geq N_{2}$, the total contribution of characters of type (II) to $\Sigma^{*}$ has absolute value $<1/6<(q+1)/16$.
Proof.
We will choose $N_{2}>4A_{2}+1$ so that $p>r$. There are fewer than $q^{2}$ choices for $s_{2p}$, at most $q^{2r}$ choices for $s_{2r}$ (up to conjugacy), and at most $q+1$ choices for the element $s_{\kappa}$. By (4.4), the condition $\chi(x)\neq 0$ implies that (after conjugating $x$ suitably) $x\in L:=\mathbf{C}_{G}(s)$ and $\psi(x)\neq 0$. Viewed as an element in $L$, the $\mathrm{GL}_{p}(q^{2})$-component of $x$ in the case of (6.2) is again regular semisimple with centralizer $C_{q^{2p}-1}$. Hence the $\mathrm{GL}_{p}(q^{2})$-component of the unipotent character $\psi$ is labeled by one of the $p$ hook partitions of $p<N/2$, and the number of possibilities for the other component of $\psi$ is bounded by the aforementioned function $f(A_{2})$. In the case of (6.3), the $\mathrm{GL}_{p+r}(q^{2})$-component of $x$ is regular semisimple with centralizer $C_{q^{2p+2r}-1}$, whence the $\mathrm{GL}_{p+r}(q^{2})$-component of the unipotent character $\psi$ is labeled by one of the $p+r$ hook partitions of $p+r\leq N/2$; also, $\psi$ is uniquely determined by this component.
Thus there are fewer than $(q+1)q^{2r+2}f(A_{2})N/2$ characters of type (II) with $\chi(x)\neq 0$. Each of them has degree
$\chi(1)\geq[G:L]_{q^{\prime}}>q^{cN^{2}}$
for some absolute constant $c>0$. Also, $|\chi(x)|\leq 4pr<2NA_{2}$ by (6.7). Hence, by (4.5), the total contribution of these characters to $|\Sigma^{*}|$ is at most
$2(q+1)q^{2r+2}f(A_{2})N^{3}A_{2}^{2}q^{-\sigma cN}<\frac{1}{6}$
when $N\geq N_{2}$ and $N_{2}$ is large enough compared to $A_{2}$. ∎
To complete the proof of (6.6) and of Theorem 5, it remains to prove:
Lemma 6.3.
For any constant $A_{2}\geq 3$, there is an absolute constant $N_{3}>0$ (depending on $A_{2}$) such that when $r\leq A_{2}$ and $N\geq N_{3}$, the total contribution of characters of types (III) and (IV) to $\Sigma^{*}$ has absolute value $<1/6<(q+1)/16$.
Proof.
We will choose $N_{3}>4A_{2}+1$ so that $p>r$. Since $\mathsf{supp}(g)\geq 2$, by Lemma 3.1 we have
$|\chi(g)|\leq\sqrt{|\mathbf{C}_{G}(g)|}\leq q^{N^{2}/2-1.7N+2.5}$
for any $\chi$; and we still have the bound (6.7) for $|\chi(x)|$.
Each $\chi$ of type (III) can be written in the form (4.3), where $L=\mathbf{C}_{G}(s)$ is given in (6.4), and
$\chi(1)\geq[G:L]_{q^{\prime}}>q^{N^{2}/2-N-c_{1}}$
for some absolute constant $c_{1}>0$ (that depends on $A_{2}$). There are fewer than $(q+1)q^{p+2r}<q^{N/2+2A_{2}}$ choices for $s$. Furthermore, there are two choices for the $\mathrm{GU}_{2}(q^{p})$-component of the unipotent character $\psi$, and at most $f(A_{2})$ choices for its other component (as in the proof of Lemma 6.3). Thus there are fewer than $2q^{N/2+2A_{2}}f(A_{2})$ choices for $\chi$ of type (III). Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$8N^{2}A_{2}^{2}q^{N/2+2A_{2}}f(A_{2})q^{-0.7N+c_{1}+2.5}\leq 8N^{2}A_{2}^{2}f(A_{2})q^{-0.2N+c_{1}+2.5+2A_{2}}<\frac{1}{12}$
when $N\geq N_{3}$ and $N_{3}$ is large enough compared to $A_{2}$.
Each $\chi$ of type (IV) can be written in the form (4.3), where $L=\mathbf{C}_{G}(s)$ is given in (6.5), and
$\chi(1)\geq[G:L]_{q^{\prime}}>q^{(N^{2}-N)/2-c_{2}}$
for some absolute constant $c_{2}>0$ (that depends on $A_{2}$). There are fewer than $(q+1)q^{N-1}<q^{N+1}$ choices for $s$. Furthermore, there is one choice for the $T_{1}$-component of the unipotent character $\psi$, and at most $f(A_{2})$ choices for its other component. Hence the total contribution of these characters to $|\Sigma^{*}|$ is at most
$4N^{2}A_{2}^{2}f(A_{2})q^{N+1}q^{-1.2N+2.5+c_{2}}\leq 4N^{2}A_{2}^{2}f(A_{2})q^{-0.2N+3.5+c_{2}}<\frac{1}{12}$
when $N\geq N_{3}$ and $N_{3}$ is large enough compared to $A_{2}$. ∎
7. Thompson’s conjecture for $\mathrm{PSU}_{N}(q)$
In this section, we prove Theorem 1. The idea is to write $N=M+L$, where $L$ is divisible by $22q+$ 2, and to find a block diagonal regular semisimple element $x\in\mathrm{SU}_{N}(q)$ of the form
(7.1) $x:=\mathrm{diag}(x_{M},x_{L}).$
We choose $x_{M}$ in such a way that $x_{M}^{\mathrm{SU}_{M}(q)}x_{M}^{\mathrm{SU}_{M}(q)}$ includes all unipotent elements and all elements of support $\geq 2$. We choose $x_{L}$ so that
(7.2) $x_{L}^{\mathrm{SU}_{L}(q)}x_{L}^{\mathrm{SU}_{L}(q)}\supset\mathbf{Z}(\mathrm{SU}_{L}(q)),$
which consists of all scalar matrices of the form $\omega I$, where $\omega\in\mu_{q+1}$. The construction of $x_{M}$ was accomplished in sections §4, §5, and §6 for many integers $M$. In this section, we determine the decompositions $N=M+L$ and the elements $x_{L}$. We start with the second task.
Let $T$ denote a finite subgroup of $\bar{\mathbb{F}}_{q}^{\times}$ which is stable under the $q$-Frobenius $\mathsf{Frob}_{q}$. There is a natural action of $T\rtimes(\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z})$ on $\bar{\mathbb{F}}_{q}^{\times}$ such that the orbit of $\alpha$ is
$\mathcal{O}_{T}(\alpha):=\{\omega\alpha^{\pm q^{i}}\mid\omega\in T,\,i\in\mathbb{N}\}.$
Slightly abusing the language, in what follows, a $T$-orbit will mean a set of the form $\mathcal{O}_{T}(\alpha)$. We say a subset of $\bar{\mathbb{F}}_{q}^{\times}$ is $T$-stable if it is stable under this action, that is, stable under $\mathsf{Frob}_{q}$, multiplication by any element of $T$, and inversion. This is equivalent to being a union of $T$-orbits $\mathcal{O}_{T}(\alpha_{i})$. We say $\alpha\in\mathbb{F}_{q^{m}}$ is $T$-regular if $\mathcal{O}_{T}(\alpha)$ has $2m|T|$ elements. In this case, it consists of $2|T|$ $\mathsf{Frob}_{q}$-orbits of length $m$, with representatives $\omega\alpha,\omega^{-1}\alpha^{-1}$, $\omega\in T$, which shows that the product of the elements in $\mathcal{O}_{T}(\alpha)$ is $1$. In particular, we say $\alpha$ is $\mathrm{SU}$-regular if it is $\mu_{q+1}$-regular.
Lemma 7.1.
The number of elements of $\mathbb{F}_{q^{m}}$ which are not $\mathrm{SU}$-regular is $O(m^{2}q^{1+m/2})$.
Proof.
If $\alpha\in\mathbb{F}_{q^{m}}$ is not $\mathrm{SU}$-regular, it must satisfy an equation of the form
(7.3) $\omega_{1}\alpha^{e_{1}q^{r_{1}}}=\omega_{2}\alpha^{e_{2}q^{r_{2}}},$
where $e_{1},e_{2}\in\{-1,1\}$, $\omega_{1},\omega_{2}\in\mu_{q+1}$, $r_{1},r_{2}\in\mathbb{Z}$, $|r_{1}-r_{2}|<m$, and $(\omega_{1},e_{1},r_{1})\neq(\omega_{2},e_{2},r_{2})$. Changing $r_{1}$ and $r_{2}$ within their residue classes (mod $m$) and exchanging them if necessary, we may assume $20\leq r_{1}\leq r_{2}\leq r_{1}+m/$ 2. Replacing $\omega_{2}$ by $\omega_{2}/\omega_{1}$, we may assume $\omega_{1}=1$. Therefore, we need only consider $O(m^{2}q)$ equations, and each has at most $q^{m/2}+1$ solutions. The lemma follows. ∎
Theorem 7.2.
Let $q\leq 7$ be a prime power and let $N$ be an integer which is not $0\pmod{2q+2}$ and is not $6$ or $7\pmod{12}$ if $q=5$. Then there exists an arbitrarily large integer $M\equiv N\pmod{2q+2}$ and an element $x\in H:=\mathrm{SU}_{M}(q)$ satisfying:
  • The element $x$ is regular semisimple,
  • The number of irreducible factors of the characteristic polynomial of $x$ as a polynomial over $\mathbb{F}_{q^{2}}$ is bounded (in fact by $5$),
  • Every unipotent element in $H$ is the product of two $H$-conjugates of $x$,
  • Every element $g\in H$ with $\mathsf{supp}(g)\geq 2$ is the product of two $H$-conjugates of $x$.
Proof.
Such elements are constructed in the proofs of Theorems 3, 4, and 5 whenever $M$ is respectively of the form $2p$, for $p$ a sufficiently large prime; $12p+$ 1, for $p$ a sufficiently large prime; or $2p+2r+\kappa$, for $p$ a sufficiently large prime, $r$ chosen from a set which represents every residue class (mod $120$) which is prime to $120$, and $\kappa\in\{0,1\}$. When $q=2$, $2p$ represents $2$ and $4$ (mod $6$), and $12p+2r+$ 1 represents $1$, $3$, and $5$. When $q=3$, $2p+2r$ represents $4$ (mod $8$), $12p+2r+$ 1 represents $1$ and $5$, $2p$ represents $2$ and $6$, and $12p+$ 1 represents $3$ and $7$. When $q=4$, $2p$ represents $2$, $4$, $6$, and $8$ (mod $10$), and $12p+2r+$ 1 represents $1$, $3$, $5$, $7$, and $9$. When $q=5$, $2p+2r$ represents $4$ and $8$ (mod $12$), $12p+2r+$ 1 represents $1$, $5$, and $9$, $2p$ represents $2$ and $10$, and $12p+$ 1 represents $3$ and $11$. When $q=7$, $2p+2r$ represents $4$, $8$, and $12$ (mod 16)16)16 ), $12p+2r+$ 1 represents $1$, $5$, $9$, and $13$, $2p$ represents $2$, $6$, $10$, and $14$, and $12p+$ 1 represents $3$, $7$, $11$, and $15$. Since $p$ can be taken arbitrarily large within each residue class (mod $22q+$ 2) which is prime to $22q+$ 2, $M$ can likewise be taken arbitrarily large, and the statement holds. ∎
We can now prove Theorem 1.
Proof.
We have already seen that the main theorem of [3] allows us to assume $q\leq 7$. If $N$ is divisible by $q+1$ and sufficiently large then $\mathrm{PSU}_{N}(q)$ satisfies Thompson’s conjecture by [15, Theorem 7.7]. This covers all cases when $N\equiv 0\pmod{2q+2}$ as well as the case that $q=5$ and $N\equiv 6\pmod{12}$. Also, by [3], if $N$ is odd and $q\geq 5$, then $\mathrm{PSU}_{N}(q)$ satisfies Thompson’s conjecture, which covers the case that $q=5$ and $N\equiv 7\pmod{12}$. Therefore, it remains to handle the case $N$ belongs to a residue class (mod $22q+$ 2) for which Theorem 7.2 applies. It follows that there exists an arbitrarily large integer $M$ congruent to $N$ (mod $22q+$ 2) and an element $x_{M}\in\mathrm{SU}_{M}(q)$ satisfying all the properties of that theorem. For each residue class, we fix $M$ and $x_{M}$.
If $N$ is sufficiently large and belongs to the same residue class as $M$, we may assume that $N=M+(2q+2)m$, where $m>M$ is large enough that Lemma 7.1 guarantees the existence of an $\mathrm{SU}$-regular element $\alpha\in\mathbb{F}_{q^{m}}$. We set $L:=(2q+2)m$. If $X\in\mathrm{SU}_{m}(q)$ has eigenvalue $\alpha$ and $\mu_{q+1}=\langle\gamma\rangle$, then
$x_{L}:=\mathrm{diag}(X,\gamma X,\gamma^{2}X,\ldots,\gamma^{q}X,X^{-1},\gamma X^{-1},\ldots,\gamma^{q}X^{-1})$
is regular semisimple. Permuting the blocks preserves the conjugacy class in $\mathrm{SU}_{L}(q)$, so (7.2) holds, and for all $\omega\in\mu_{q+1}$, the scalar matrix $\omega I$ is a product of two conjugates of $x_{L}$.
The irreducible factors of the characteristic polynomial of $x_{M}$ have degree $\leq M<m$, so it cannot have an irreducible factor in common with the characteristic polynomial of $x_{L}$. Therefore, the element $x$ with blocks $x_{M}$ and $x_{L}$ defined in (7.1) is regular semisimple, and its characteristic polynomial has a bounded number of factors, independent of all choices. By [14, Theorem 1.1], there exists $B>0$, independent of all choices, such that every element $g\in\mathrm{SU}_{N}(q)$ with $\mathsf{supp}(g)>B$ lies in $x^{\mathrm{SU}_{N}(q)}x^{\mathrm{SU}_{N}(q)}$. As $B$ is independent of choices, we may assume $M>B$.
Defining $C$ to be the conjugacy class of the image of $x$ in $S$, we claim that $C^{2}=\mathrm{PSU}_{N}(q)$. To prove this claim, suppose we have any element of $\mathrm{PSU}_{N}(q)$. We lift it to $g\in\mathrm{SU}_{N}(q)$ in such a way that if $g$ is a scalar multiple of a unipotent element, then the scalar is $1$. We may assume that $\mathsf{supp}(g)\leq B<M<L$, so $g$ has a (unique) primary eigenvalue $\lambda\in\mu_{q+1}$ whose eigenspace has dimension $\geq N-B>L$. It follows that $g$ is conjugate to a block diagonal matrix $(g_{M},g_{L})\in\mathrm{SU}_{M}(q)\times\mathrm{SU}_{L}(q)<\mathrm{SU}_{N}(q)$, where $g_{L}=\lambda I$ and $\lambda$ is an eigenvalue of $g_{M}$. Hence, if $g_{M}$ is a scalar multiple of a unipotent element, the scalar must be $\lambda$, so $g$ itself is $\lambda$ times a unipotent element, and so, by our choice of $g$, the scalar must be $1$.
By Theorem 7.2 (and [9, Lemma 5.1] if $g_{M}$ is semisimple of support $1$), $g_{M}$ is the product of two $\mathrm{SU}_{M}(q)$-conjugates of $x_{M}$. By construction, $g_{L}$ is the product of two $\mathrm{SU}_{L}(q)$-conjugates of $x_{L}$, so we are done. ∎